### v1

# Near-Optimal Adaptive Transforms for LLM Quantization

Abstract
Quantization to low bitwidth is a standard approach for deploying large language models, however, a few extreme weights and activations stretch the dynamic range and reduce the effective resolution of the quantizer. A common mitigation approach is to apply some fixed orthogonal transforms, such as Hadamard matrices, before quantization, which typically reduces the dynamic range. Yet, these transforms ignore the statistics of the data, and their optimality is currently not understood. In this work, we derive, for the first time, closed-form optimal linear blockwise transforms for joint weight-activation quantization using standard data-free quantizers for common numerical formats. Specifically, we provide derivations of the optimal adaptive (data-aware) transforms for round-to-nearest (RTN), AbsMax-scaled block quantizers for both integer and floating-point formats. The resulting construction, which we call WUSH, combines a Hadamard backbone with a data-dependent component based on second-order moments, yielding a non-orthogonal transform that is provably optimal under mild assumptions and remains structured for efficient implementation. Preliminary experimental results show that our approach consistently improves upon the Hadamard transform for common formats.
1 Introduction
1.1 Overview
Quantization of model weights (NEURIPS2022_c3ba4962; frantar2023optq; MLSYS2024_42a452cb) or activations (pmlr-v202-xiao23c; NEURIPS2024_b5b93943) is now a standard tool for shrinking and accelerating large language models (LLMs) (kurtic2025give), making low-precision inference feasible on a wide range of hardware. A central difficulty, however, is that a few extreme “outlier” weights and activations expand the dynamic range and thereby degrade the effective resolution of low-bit representations.
It is known that one way to mitigate these outliers is to apply linear transforms before quantization (jegou2008hamming; jegou2010product); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (cheequip2023; pmlr-v235-tseng24a; NEURIPS2024_b5b93943; liu2025spinquant). For LLMs, Hadamard rotations have been remarkably effective, and blockwise variants aligned with quantization groups have also been shown to be useful in practice. The recent work of egiazarian2025bridginggappromiseperformance shows that the block-wise Hadamard transform offers the best empirical performance for quantization among a large set of existing orthogonal transforms. Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not depend on the statistics of the underlying weights or activations. This raises a natural question: if the Hadamard transform is not data-aware, in what sense can it be considered optimal for quantization?
In this work, we address this question by deriving closed-form optimal linear blockwise transforms for joint weight-activation quantization, which are generally non-orthogonal and adaptive, i.e., data-aware. As opposed to prior methods such as SpinQuant (liu2025spinquant) or FlatQuant (sun2024flatquant), our approach is closed-form and calibration-based, as it does not require training or fine-tuning.
We call our construction WUSH, and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers. The name WUSH is a mnemonic for its composition. For a layer with blockwise weight $\bm{W}_{\left(i\right)}$ and calibration activation $\bm{X}_{\left(i\right)}$ in the $i$-th block, the WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}$ is constructed as
$\bm{T}_{\mathrm{hsuw}}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
where $\bm{W}_{\left(i\right)}^{\prime}$ is the Cholesky decomposition of $\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}=\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}$, the matrices $\bm{U}_{\left(i\right)},\bm{S}_{\left(i\right)},\bm{V}_{\left(i\right)}$ are from the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$, and $\bm{H}$ is a Hadamard matrix. We will define it again in a more exhaustive manner in Eq. (6). WUSH is explicitly adapted to minimize loss in second-order statistics while retaining a structure amenable to efficient implementation and providing a principled, data-aware alternative to heuristic Hadamard-style and blockwise rotations.
1.2 Related Work
Post-training quantization and outliers. Generic post-training quantization schemes, such as RTN (round-to-nearest), OBQ (frantar2022obc), and GPTQ (frantar2023optq), are highly sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values determine the quantization scale. As a result, most quantization levels are spent covering rare extremes instead of the bulk of the distribution, which leads to a significant loss of accuracy.
Non-uniform bitwidths and explicit outlier storage. One common strategy to mitigate outliers is to use non-uniform bitwidths. Methods such as LLM.int8() (NEURIPS2022_c3ba4962), SpQR (dettmers2024spqr), and QUIK (ashkboos-etal-2024-quik) explicitly separate and store outliers in higher precision, while HPTQ (chen2025geometryllmquantizationgptq) uses Huffman encoding to compress the outliers and reduce their storage costs. These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels and additional engineering effort for efficient deployment.
Transform-based outlier mitigation. Another line of work applies transforms to the weights and activations before quantization to reduce the impact of outliers. SmoothQuant (pmlr-v202-xiao23c) and AWQ (MLSYS2024_42a452cb) rescale channels to stabilize the dynamic ranges of weights and activations. QuIP (cheequip2023) introduces an incoherence processing step. QuIP# (pmlr-v235-tseng24a), QTIP (NEURIPS2024_6de2e84b), and QuaRot (NEURIPS2024_b5b93943) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (liu2025spinquant) learns rotations optimized on calibration data. However, their transforms are either heuristic, costly to learn, or primarily designed for weight-only quantization, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
Blockwise transforms and emerging FP formats. The recent FP formats, such as MXFP (ocp_mx_2023) and NVFP (nvidia_blackwell_2024), have motivated blockwise transform schemes. shao2025blockrotationneedmxfp4 indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but egiazarian2025bridginggappromiseperformance highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions. In contrast, our work derives a closed-form, data-aware optimal blockwise transform (WUSH) and analyzes how such a transform interacts with both FP and INT block quantizers.
1.3 Methodology
Problem setup. We formulate the transformed weight-activation quantization problem as follows. Define $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ as the weight of a linear layer with $d_{\mathrm{in}}$ input channels and $d_{\mathrm{out}}$ output channels. Define $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$ as the calibration input activation with $d_{\mathrm{in}}$ embedding channels and $d_{\mathrm{batch}}$ tokens. Define $q\left(\cdot\right)$ as a quantizer that maps a matrix of continuous values to quantized values (we will specify $q$ later). Let $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}$ be the transforms applied to $\bm{W}$ and $\bm{X}$, respectively. For weight-activation quantization, the output activation $\bm{W}^{\top}\bm{X}$ becomes $q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)$. Our goal is to choose $\bm{T}_{\mathrm{W}}$, $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
$\ell=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)-\bm{W}^{\top}\bm{X}\right\|_{\mathrm{F}}^{2}$ (2)
is minimized under the given quantizer $q$. The constant $d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}$ is added to simplify the analysis later.
Block-independent constraint. While the weights can be pre-quantized, the activations must be transformed and quantized dynamically at inference time. To reduce the computational overhead and simplify the problem, we constrain the transforms $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ to be block-diagonal, and the quantizer $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales. We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$ ($d$ is a factor of $d_{\mathrm{in}}$, and $d$ is a power of $2$). Denote
$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right)\quad\text{and}\quad\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right)$ (3)
with $\bm{T}_{\mathrm{W}_{\left(i\right)}},\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathrm{R}^{d\times d}$. We partition $\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ and $\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{X}_{\left(i\right)}\in\mathrm{R}^{d\times d_{\mathrm{batch}}}$. Then, we can express the output without quantization as $\bm{W}^{\top}\bm{X}=\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}$ and the output with weight-activation quantization as
$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$ (4)
Finally, define the blockwise output loss as
$\ell_{\left(i\right)}=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right)-\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}\right\|_{\mathrm{F}}^{2}.$ (5)
We approximate the layerwise loss as $\ell\approx\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise loss $\ell_{\left(i\right)}$ independently.
Optimal block-diagonal transforms. We will prove that there exist closed-form optimal transforms for each block under reasonable assumptions and smooth approximations, which are optimal for FP and asymptotically optimal for INT quantizations. Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy $\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$, respectively. Without loss of generality (any differences will be canceled in later steps), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$. In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moment before the Cholesky decomposition. Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$. Let $\bm{H}\in\mathbb{R}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix where each element is $\pm d^{-\frac{1}{2}}$. The optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ are constructed as
${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top}\quad\text{and}\quad{\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$ (6)
respectively. Note that ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}={\bm{T}_{\mathrm{hsuw}}^{-\top}}_{\left(i\right)}$  [^1] The -$\top$ superscript notation means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$. , which can be easily verified by calculating ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}{\bm{T}_{\mathrm{hsuw}}^{\top}}_{\left(i\right)}=\mathbf{I}$ with $\mathbf{I}$ as the identity matrix.
Remark. The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Eq. (6), which explains why it has been empirically shown to be one of the most effective data-agnostic orthogonal transforms.
2 Theoretical Derivation
In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$. We omit the subscript $\left(i\right)$ to simplify notation.
2.1 Problem Setup and General Proof Approach
Probabilistic reformulation. We reformulate the problem from a probabilistic perspective. We split the matrices $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{d_{\mathrm{out}}}\right]$ and $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{d_{\mathrm{batch}}}\right]$ into columns $\bm{w}_{k},\bm{x}_{k}\in\mathbb{R}^{d}$ and treat the columns as i.i.d. samples from $d$-dimensional independent distributions $\bm{w}\sim\mathcal{D}_{\mathrm{w}}$ and $\bm{x}\sim\mathcal{D}_{\mathrm{x}}$, respectively. Then the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy $\bm{W}^{\prime}\bm{W}^{\prime\top}=\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top}$ and $\bm{X}^{\prime}\bm{X}^{\prime\top}=\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}$, respectively. We can reinterpret the blockwise loss as $\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}$.
Unbiased quantization. We assume that $q$ is a stochastic and unbiased quantizer, i.e., for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, the quantization error $\varepsilon\left(\bm{\alpha}\right)=q\left(\bm{\alpha}\right)-\bm{\alpha}$ is a random vector with $\mathbb{E}_{\varepsilon\left(\bm{\alpha}\right)}\varepsilon\left(\bm{\alpha}\right)=\bm{0}$. We further constrain $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$ such that $q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)$ is unbiased with respect to $\bm{w}^{\top}\bm{x}$. Then, using the first-order approximation, we can split $\ell$ into two non-negative terms:
$\ell=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\left(\bm{T}_{\mathrm{W}}\bm{w}+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{\top}\left(\bm{T}_{\mathrm{X}}\bm{x}+\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)-\bm{w}^{\top}\bm{x}\right)^{2}$ (7)
$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
$\approx$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right)$
$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$
Problem reduction. We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify if $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$. The two terms have similar forms and can be optimized in similar ways. We focus on finding the optimal $\bm{T}_{\mathrm{X}}$ for the second term. Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}\sim\mathcal{D}_{\mathrm{y}}$. Define the transformation matrix $\bm{T}=\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}$. Then, the second term becomes
$\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\left(\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}\right)^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\left(\bm{W}^{\prime-\top}\bm{y}\right)\right)\right\|^{2}$ (8)
$=$ $\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$
Therefore, the two-sided quantization problem is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute $\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}$.
Transformation parameterization. Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$; thus, $\bm{T}$ can be parameterized as $\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}$. Denote $\bm{S}=\operatorname{diag}\left(s_{1},\dots,s_{d}\right)$ and $\bm{S}^{\prime}=\operatorname{diag}\left(s^{\prime}_{1},\dots,s^{\prime}_{d}\right)$. Without loss of generality, assume $s_{1}\geq\dots\geq s_{d}>0$ and $s^{\prime}_{1}\geq\dots\geq s^{\prime}_{d}>0$.
Theorem 1 (Optimal Transform)
The optimal configuration of $\bm{T}$ for floating-point (FP) data types is $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$. For integer (INT), the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
Proof We provide detailed proofs for FP and INT in Sections˜2.3 and 2.2, respectively. For each data type, we apply a two-step proof strategy. We first provide a smooth modeling of the quantizer and treat the type casting errors as random noise. Secondly, we calculate the expected quantization error with respect to the parameterized transform and minimize this expectation by solving the corresponding optimization problem.
Some useful notations and basic results are as follows:
  • Denote the $k$-th standard basis vector as $\mathbf{e}_{k}\in\left\{0,1\right\}^{d}$, where the $k$-th element is $1$ and 0 otherwise.
  • The second moment of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}=\bm{W}^{\prime\top}\left(\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}\right)\bm{W}^{\prime}=\bm{W}^{\prime\top}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{W}^{\prime}=\bm{U}\bm{S}^{2}\bm{U}^{\top}$.
  • The second moment of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}=\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}=\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}$.
  • The expected squared norm of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$.
  • Similarly, the expected squared norm of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$.
  • Because $\bm{S}\succ\bm{0}$, $\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\right)$. By Jensen’s inequality, $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$.
2.2 Proof for Floating-Point (FP) Types
2.2.1 Quantization Error Modeling
Theorem 2 (FP Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$. Assume we do not need to consider the subnormal and special number cases. The tuple $\left(s,e,m\right)$ represents the number $x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right)$.
Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$, we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as $x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ with $2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ being a $a+b$ bit unsigned integer (concatenating the binary representations of $e$ and $m$). The minimum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ is $\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{-2^{a-1}+1}$ and the maximum is $\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1}$.
Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type. Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely. $\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in the logarithm space and first-order expansion. For $x\neq 0$,
$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}=\mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$ (9)
$\approx$ $\ \mathrm{sgn}\left(x\right)\left(2^{2^{-b}2^{b}\log_{2}\left|x\right|}+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{2^{-b}2^{b}\log_{2}\left|x\right|}2^{-b}\ln 2\right)$
$=$ $\ x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right).$
For $x=0$, $\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0$. We model the casting error $\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
[…]
$q\left(x\right)=$ $\ \mathrm{SE}a\mathrm{M}b\left(xs^{-1}\right)\mathrm{SE}\tilde{a}\mathrm{M}\tilde{b}\left(s\right)=\left(1+\left(\ln 2\right)2^{-b}\xi\right)xs^{-1}\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)s$ (10)
$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)x$
with the independent random variables $\xi,\tilde{\xi}\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
Define $\eta=\left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)-1$. The quantization error is $q\left(x\right)-x=\eta x$.
$\mathbb{E}_{\eta}\eta=\mathbb{E}_{\xi,\tilde{\xi}}\left(\ln 2\right)2^{-b}\xi+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}+\left(\ln 2\right)^{2}2^{-b-\tilde{b}}\xi\tilde{\xi}=0.$ (11)
2.2.2 Loss Minimization
Using Theorem˜2, the minimization objective is
$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (12)
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\operatorname{tr}\left(\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right)$
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\right)\bm{T}^{-\top}\right)$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
where $\odot$ represents elementwise multiplication. The lower bound of the trace term in the loss is
$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (13)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)\bm{T}^{-\top}\right)$
$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (rotation-invariance of $L_{2}$ norm)
$\geq$ $\ \sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Cauchy-Schwarz)
$=$ $\ d\left(d^{-1}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}\right)$
$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen)
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\right)\right)^{2}$
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$
The equality can be attained by choosing $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$ such that
$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (14)
$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$
2.2.3 Discussion
Under our smooth modeling in Theorem˜2, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$ (15)
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right),$
so the orthogonal transforms will not be helpful for reducing the quantization error. Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error by at most $d$ times, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario. Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term $\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$, which is the same as the case of orthogonal $\bm{T}$.
2.3 Proof for Integer (INT) Data Types
2.3.1 Quantization Error Modeling
We model the quantization error using pseudo noise similar to DiffQ (defossez2022differentiable).
Theorem 3 (INT Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof Define the scalar quantizer of $b$-bit symmetric integer as $q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s$, where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$. To simplify the analysis, we model the quantizer as $q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$. Define $\eta=2^{1-b}\xi$, then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$. The quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$.
2.3.2 Loss Minimization
Using Theorem˜3, the minimization objective is
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (16)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}.$
We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$. Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality, we obtain:
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)\geq\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right).$ (17)
Equality is attained when $\bm{R}=\mathbf{I}$.
Next, we obtain the lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$. Express the expectations as $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$. By the relationships among the mean, maximum, and summation values, we have:
$d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq d\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},$ (18)
where $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ and $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$.
Lemma 4 (Maximum Inequalities for Tail-Bounded Distributions)
If $X_{1},\dots,X_{d}\in\mathbb{R}$ are zero-mean Gaussian random variables,
$\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}.$ (19)
If $X_{1},\dots,X_{d}\in\mathbb{R}$ are zero-mean Laplacian random variables,
$\mathbb{E}\max_{k}X_{k}^{2}\leq\left(2^{-1}\left(\ln d\right)^{2}+\ln d+1\right)\max_{k}\mathbb{E}X_{k}^{2}.$ (20)
Both inequalities hold without the need for independence.
Proof We provide proofs for Gaussian and Laplacian distributions in Sections˜A.1 and A.2, respectively.
Using Lemma 4, for a tail-bounded $\mathcal{D}_{\mathrm{y}}$ that is a zero-mean multivariate Gaussian or Laplacian, we can further tighten the upper bound by
$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\min\left\{d^{o\left(1\right)}\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\right\}.$ (21)
When $\bm{U}^{\prime}=\bm{H}$, the marginal second moment becomes $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)=d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$, which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$. Taken together, by setting $\bm{R}=\mathbf{I}$ and $\bm{U}^{\prime}=\bm{H}$, the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ is bounded by $\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ within a gap of $d$ such that
$d^{-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (22)
And for a tail-bounded $\mathcal{D}_{\mathrm{y}}$, the gap is narrowed to $d^{o\left(1\right)}$ such that
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (23)
By the Cauchy-Schwarz inequality,
$\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\geq\left(\operatorname{tr}\left(\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)^{\frac{1}{2}}\left(\bm{S}^{\prime 2}\right)^{\frac{1}{2}}\right)\right)^{2}=\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (24)
and the equality is attained when $\bm{S}^{\prime}\propto\bm{S}^{\frac{1}{2}}$. Without loss of generality, we can choose $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ to minimize both the lower and upper bounds of the quantization loss.
2.3.3 Discussion
Consider the case of $\bm{T}$ being orthogonal. $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$. For any distribution $\mathcal{D}_{\mathrm{y}}$, the bounds are $\operatorname{tr}\left(\bm{S}^{2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq d\operatorname{tr}\left(\bm{S}^{2}\right)$. For the Hadamard $\bm{T}=\bm{H}$ and a tail-bounded $\mathcal{D}_{\mathrm{y}}$, $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)}\operatorname{tr}\left(\bm{S}^{2}\right)$ with high probability. Because $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$, an orthogonal $\bm{T}$ is suboptimal, and the error bounds of an orthogonal $\bm{T}$ can be at most $d$ times larger than those of the optimal $\bm{T}$, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario.
3 Experiments
3.1 Algorithm
1
Input: weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, damping ratio $\lambda\in\mathbb{R}$
Output: WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\hat{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
2
[…]
3
[…]
5for $i\leftarrow 1\ \mathrm{to}\ d_{\mathrm{in}}/d\ \mathrm{(in\ parallel)}$ do
6
7 $j_{1},j_{2}\leftarrow\left(i-1\right)d+1,id$
[…]
8
[…]
9
 $\bm{W}^{\prime}\leftarrow\operatorname{Cholesky}\left(\bm{M}_{\mathrm{W}}\right)$ // $\bm{W}^{\prime}\bm{W}^{\prime\top}=\bm{M}_{\mathrm{W}}$
10
 $\bm{U},\bm{\Lambda}\leftarrow\operatorname{Eigendecomposition}\left(\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}\right)$ // $\bm{U}\bm{\Lambda}\bm{U}^{\top}=\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}$
11
 ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\leftarrow\bm{H}\bm{\Lambda}^{-\frac{1}{4}}\bm{U}^{\top}\bm{W}^{\prime\top}$ // WUSH transform
12
 $\hat{\bm{W}}_{\left(i\right)}=q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}^{-\top}\bm{W}[j_{1}:j_{2},1:d_{\mathrm{out}}]\right)$ // pre-quantize weight block
13
14 end for
15
Algorithm 1 Compute WUSH Transforms and Pre-Quantize Weights
Algorithm˜1 summarizes the procedure used to compute the blockwise transforms and pre-quantized weights for each linear layer. In practice, the required second-order information can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ. As a result, the calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one.
During inference, the forward pass is calculated as
$\sum_{i=1}^{d_{\mathrm{in}}/d}\hat{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\bm{X}_{\left(i\right)}\right).$ (25)
with the symbol $\bm{X}$ being overloaded as the new activations instead of the calibration ones.
3.2 Layer Loss
Table 1: Weight-activation RTN quantization losses with identity (I), random rotation (R), Hadamard (H), WUSH without Hadamard (WUS), and WUSH transforms.
Format Transform Q K V O Gate Up Down
MXFP4 I .01105 .01201 .01068 .00435 .00710 .00656 .00547
R .00761 .00773 .00914 .00384 .00556 .00568 .00407
H .00724 .00720 .00860 .00379 .00545 .00561 .00390
WUS .00627 .00722 .00405 .00357 .00576 .00475 .00446
WUSH .00334 .00334 .00330 .00276 .00449 .00439 .00339
NVFP4 I .00423 .00435 .00437 .00234 .00349 .00341 .00241
R .00498 .00501 .00586 .00254 .00363 .00368 .00266
H .00560 .00562 .00670 .00258 .00371 .00379 .00278
WUS .00226 .00236 .00230 .00200 .00304 .00301 .00223
WUSH .00240 .00244 .00228 .00192 .00309 .00302 .00233
INT4 I .17004 .12277 .01315 .00455 .05585 .00983 .01934
R .00579 .00584 .00696 .00294 .00422 .00429 .00310
H .00557 .00555 .00680 .00286 .00409 .00425 .00303
WUS .21268 .14233 .01067 .00454 .05021 .00742 .01308
WUSH .00239 .00243 .00254 .00210 .00343 .00343 .00255
We report the layerwise weight-activation RTN quantization loss ($L_{2}$ loss normalized by the number of elements) for each linear layer in the 18th transformer block of Qwen3-8B, using 32 calibration samples of sequence length 2048 from the FineWeb-Edu (NEURIPS2024_370df50c) dataset. We compare the WUSH transform with the identity (I), random rotation (R) averaged over 10 runs, Hadamard (H), and WUSH without Hadamard (WUS). Table˜1 summarizes the losses for MXFP4, NVFP4, and INT4 formats, where INT4 denotes 4-bit integers with Gaussian MSE clipping and BF16 scales of group size 32. Across formats, WUSH substantially reduces the loss: WUSH consistently yields the smallest loss for MXFP4 and INT4, while WUSH and WUS are almost equally optimal for NVFP4.
These empirical trends are broadly consistent with our theoretical analysis, suggesting that the approximation assumptions (Section˜2.1) and the smooth error models (Sections˜2.2.1 and 2.3.1) capture the dominant sources of quantization loss in practice. At the same time, MXFP4 behaves as a hybrid between ideal FP and INT quantization: the effective mantissa step changes uniformly, imparting INT-like behavior, especially in the subnormal regime and small exponent ranges. This helps explain why Hadamard transforms still provide measurable gains for the MXFP4 format, despite the ideal FP theory (Section˜2.2.3) predicting no benefit for purely orthogonal transforms.
3.3 LLM Benchmarks
We now present preliminary experiments evaluating the accuracy of models using the WUSH transform, relative to the identity (standard RTN) and Hadamard. We conduct our experiment on the LLama-3.2-3B-Instruct, LLama3.1-8B-Instruct, Qwen3-8B, and Qwen3-14B models. We use Platinumbench (vendrow2025large) and the LM Eval Harness (gao2021framework) for evaluation.
The full results for LM Eval Harness tasks are shown in Table˜2. Here, we observe that WUSH achieves the highest average recovery score among all the methods tested, often by large margins. Remarkably, on large models such as Qwen3-14B, it helps narrow the accuracy gap between MXFP4 and NVFP4 to nearly 0.5%.
We plot the results for Platinumbench in Figures˜3, 4, 1 and 2. Again, we observe that WUSH provides significant accuracy improvements, especially in the case of MXFP.
Table 2: Accuracy results on the LM Eval Harness for round-to-nearest (RTN) quantization of weights and activations with identity (I), Hadamard (H), and WUSH transforms.
Model Format Transform MMLU GSM8K HellaSwag WinoGrande Avg. Recovery%
Llama-3.2-3B-Instruct BF16 - 64.43 78.01 73.42 70.09 71.49
NVFP4 I 61.02 71.04 70.90 66.77 67.43 94.33
H 59.91 64.82 69.77 65.59 65.02 90.95
WUSH 61.25 71.78 71.08 67.18 67.82 94.87
MXFP4 I 56.81 60.80 67.30 64.56 62.37 87.25
H 55.99 61.11 68.36 64.09 62.39 87.27
WUSH 58.45 68.60 69.99 66.54 65.89 92.17
Llama-3.1-8B-Instruct BF16 - 72.76 85.06 80.01 77.90 78.93
NVFP4 I 68.64 80.29 77.97 74.43 75.33 95.44
H 66.96 77.41 77.29 74.35 74.00 93.75
WUSH 68.83 78.57 78.22 75.47 75.28 95.37
MXFP4 I 62.21 69.52 73.76 72.61 69.53 88.08
H 63.03 75.44 75.51 71.90 71.47 90.54
WUSH 66.85 75.16 77.28 73.56 73.21 92.75
Qwen3-8B BF16 - 72.98 90.90 75.52 70.56 77.49
NVFP4 I 70.47 88.40 74.44 69.53 75.71 97.70
H 70.19 86.35 73.02 68.11 74.42 96.03
WUSH 70.86 89.33 74.82 69.79 76.20 98.33
MXFP4 I 67.69 84.23 71.24 67.40 72.64 93.74
H 67.53 87.04 71.48 67.64 73.42 94.75
WUSH 70.29 87.95 74.06 69.50 75.45 97.36
Qwen3-14B BF16 - 77.18 91.96 79.84 74.27 80.81
NVFP4 I 75.20 90.22 77.38 73.32 79.03 97.80
H 74.98 92.04 77.76 72.38 79.29 98.12
WUSH 75.43 91.78 78.82 73.18 79.80 98.75
MXFP4 I 72.92 90.22 76.68 71.51 77.83 96.31
H 73.13 87.41 76.29 71.67 77.13 95.44
WUSH 74.56 91.87 78.18 73.15 79.44 98.30
Figure 1: Comparison of different transforms on Qwen3-8B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 2: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 3: Comparison of different transforms on Llama3.2-3B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 4: Comparison of different transforms on Llama3.1-8B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
4 Future Work
A natural next step is to integrate WUSH with advanced post-training quantization schemes such as GPTQ. In this setting, the weights are updated iteratively through error propagation, while WUSH is constructed from the second-order statistics of the updated weights. Designing an algorithm that jointly updates the weights and the associated blockwise transforms is non-trivial and will require adapting both the optimization schedule and the calibration pipeline.
From a systems perspective, efficient kernel support for WUSH is another important direction. Unlike blockwise Hadamard rotations, which reuse a single transform across all blocks, WUSH generally assigns a distinct transform to each block. This per-block specialization complicates kernel reuse and fusion in existing low-level libraries.
Finally, it will be useful to explore structurally simplified variants of WUSH that trade optimality for lower computational overhead. One promising avenue is to approximate the data-aware component with a diagonal matrix, thereby reducing the cost of applying the transforms at inference time.
5 Conclusion
We have studied blockwise weight-activation quantization for LLMs in the presence of heavy-tailed outliers and have shown that, under mild assumptions, this problem admits closed-form linear transforms for both FP and INT block quantizers. Our construction, WUSH, combines a fixed Hadamard backbone with a data-aware component derived from second-order statistics, yielding a non-orthogonal blockwise transform that is provably optimal and remains amenable to efficient implementation. These results clarify the empirical success of Hadamard rotations and indicate that principled, data-aware block transforms can significantly narrow the gap between low-bit and full-precision models.
Acknowledgment
We would like to thank Tijmen Blankevoort for useful discussions and for suggesting the name WUSH, which is a clear improvement over our previous name (HSUW).
Appendix A Maximum Inequalities for Tail-Bounded Distributions
Lemma 5 (Expectation from the Survival Function)
Let $X\geq 0$ be a random variable with the density function $f$ and the survival function $S(t)=\mathbb{P}\left(X\geq t\right)$ with $t\geq 0$. Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define $G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u$. Assume that $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$. Then $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
Proof Since $X$ has density $f$ and takes values in $0,\infty)$, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t$. The survival function satisfies $S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f(u)\mathrm{d}u$, so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$. Hence, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t$. We now integrate by parts, using $G^{\prime}\left(t\right)=g\left(t\right)$,
$-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t=-\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}+\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (26)
By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$. Therefore, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
A.1 Gaussian Distribution
Definition of sub-Gaussian.
A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if $\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
Basic properties of sub-Gaussian distributions.
The variance proxy is larger than or equal to the variance: $\sigma^{2}\geq\mathbb{E}\left(X-\mathbb{E}X\right)^{2}$. For a Gaussian distribution, the optimal variance proxy is the same as the variance.
Chernoff bound on the tail: $\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ for all $t\geq 0$.
Maximum inequality.
Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$. Denote $\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ as the smallest variance proxy for $X_{1},\dots,X_{d}$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (27)
$\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
$=$ $\ t_{\star}^{2}+4d\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t_{\star}^{2}\right).$
Splitting at $t_{\star}=\sigma_{\star}\sqrt{2\ln\left(2d\right)}$ gives
$\mathbb{E}\max_{k}X_{k}^{2}\leq 2\sigma_{\star}^{2}\ln\left(2d\right)+2\sigma_{\star}^{2}=\left(2\ln\left(2d\right)+2\right)\max_{k}\mathbb{E}X_{k}^{2}$ (28)
without needing the independence of $X_{k}$. Also, considering the trivial bound $\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2}$, then $\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}$.
A.2 Laplacian Distribution
For a random variable $X\sim\mathrm{Laplace}\left(\mu,b\right)$ and $t\geq 0$,
$\mathbb{P}\left(\left|X-\mu\right|\geq t\right)=2\int_{t}^{\infty}2^{-1}b^{-1}\exp\left(-b^{-1}x\right)\mathrm{d}x=\exp\left(-b^{-1}t\right).$ (29)
The variance $\mathbb{E}\left(X-\mathbb{E}X\right)^{2}=2b^{2}$.
Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$. Denote $b_{\star}=\max_{k}b_{k}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}$. For $t\geq 0$, $\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right)$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (30)
$\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
$=$ $\ t_{\star}^{2}+2db_{\star}\left(t_{\star}+b_{\star}\right)\exp\left(-b_{\star}^{-1}t_{\star}\right).$
Splitting at $t_{\star}=b\ln d$ gives
$\mathbb{E}\max_{k}X_{k}^{2}\leq b_{\star}^{2}\left(\left(\ln d\right)^{2}+2\ln d+2\right)=\left(2^{-1}\left(\ln d\right)^{2}+\ln d+1\right)\max_{k}\mathbb{E}X_{k}^{2}$ (31)
without needing the independence of $X_{k}$.

### v2

@@
-Near-Optimal Adaptive Transforms for LLM Quantization
+WUSH: Near-Optimal Adaptive Transforms for LLM Quantization
 Abstract
-Quantization to low bitwidth is a standard approach for deploying large language models, however, a few extreme weights and activations stretch the dynamic range and reduce the effective resolution of the quantizer.
-A common mitigation approach is to apply some fixed orthogonal transforms, such as Hadamard matrices, before quantization, which typically reduces the dynamic range.
-Yet, these transforms ignore the statistics of the data, and their optimality is currently not understood.
-In this work, we derive, for the first time, closed-form optimal linear blockwise transforms for joint weight-activation quantization using standard data-free quantizers for common numerical formats.
-Specifically, we provide derivations of the optimal adaptive (data-aware) transforms for round-to-nearest (RTN), AbsMax-scaled block quantizers for both integer and floating-point formats.
-The resulting construction, which we call WUSH, combines a Hadamard backbone with a data-dependent component based on second-order moments, yielding a non-orthogonal transform that is provably optimal under mild assumptions and remains structured for efficient implementation.
-Preliminary experimental results show that our approach consistently improves upon the Hadamard transform for common formats.
+[…]
 1 Introduction
-1.1 Overview
-Quantization of model weights (NEURIPS2022_c3ba4962; frantar2023optq; MLSYS2024_42a452cb) or activations (pmlr-v202-xiao23c; NEURIPS2024_b5b93943) is now a standard tool for shrinking and accelerating large language models (LLMs) (kurtic2025give), making low-precision inference feasible on a wide range of hardware.
+Quantization of model weights (Dettmers et al., 2022; Frantar et al., 2023; Lin et al., 2024) or activations (Xiao et al., 2023; Ashkboos et al., 2024b) is now a standard tool for shrinking and accelerating large language models (LLMs) (Kurtic et al., 2025), making low-precision inference feasible on a wide range of hardware.
 A central difficulty, however, is that a few extreme “outlier” weights and activations expand the dynamic range and thereby degrade the effective resolution of low-bit representations.
-It is known that one way to mitigate these outliers is to apply linear transforms before quantization (jegou2008hamming; jegou2010product); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (cheequip2023; pmlr-v235-tseng24a; NEURIPS2024_b5b93943; liu2025spinquant).
+One way to mitigate such outliers is to apply linear transforms before quantization (Jegou et al., 2008; Jégou et al., 2011); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (Chee et al., 2023; Tseng et al., 2024a; Ashkboos et al., 2024b; Liu et al., 2025).
 For LLMs, Hadamard rotations have been remarkably effective, and blockwise variants aligned with quantization groups have also been shown to be useful in practice.
-The recent work of egiazarian2025bridginggappromiseperformance shows that the block-wise Hadamard transform offers the best empirical performance for quantization among a large set of existing orthogonal transforms.
-Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not depend on the statistics of the underlying weights or activations.
+Egiazarian et al.
+(2025) recently showed that the blockwise Hadamard transform offers the best empirical performance for quantization among a large set of existing transforms.
+Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not adapt to the statistics of the underlying weights or activations.
 This raises a natural question: if the Hadamard transform is not data-aware, in what sense can it be considered optimal for quantization?
 In this work, we address this question by deriving closed-form optimal linear blockwise transforms for joint weight-activation quantization, which are generally non-orthogonal and adaptive, i.e., data-aware.
-As opposed to prior methods such as SpinQuant (liu2025spinquant) or FlatQuant (sun2024flatquant), our approach is closed-form and calibration-based, as it does not require training or fine-tuning.
-We call our construction WUSH, and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers.
-The name WUSH is a mnemonic for its composition.
-For a layer with blockwise weight $\bm{W}_{\left(i\right)}$ and calibration activation $\bm{X}_{\left(i\right)}$ in the $i$-th block, the WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}$ is constructed as
-$\bm{T}_{\mathrm{hsuw}}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
+As opposed to prior methods such as SpinQuant (Liu et al., 2025) or FlatQuant (Sun et al., 2025), our approach is closed-form and calibration-based, as it does not require training or fine-tuning the entire model.
+We call our construction WUSH and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers.
+(The name WUSH is a mnemonic for its composition.) Consider a block of layer weights $\bm{W}_{\left(i\right)}$, with calibration activation $\bm{X}_{\left(i\right)}$.
+Then, the WUSH transform ${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}$ is:
+${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
 where $\bm{W}_{\left(i\right)}^{\prime}$ is the Cholesky decomposition of $\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}=\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}$, the matrices $\bm{U}_{\left(i\right)},\bm{S}_{\left(i\right)},\bm{V}_{\left(i\right)}$ are from the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$, and $\bm{H}$ is a Hadamard matrix.
-We will define it again in a more exhaustive manner in Eq.
-(6).
-WUSH is explicitly adapted to minimize loss in second-order statistics while retaining a structure amenable to efficient implementation and providing a principled, data-aware alternative to heuristic Hadamard-style and blockwise rotations.
-1.2 Related Work
-Post-training quantization and outliers.
-Generic post-training quantization schemes, such as RTN (round-to-nearest), OBQ (frantar2022obc), and GPTQ (frantar2023optq), are highly sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values determine the quantization scale.
-As a result, most quantization levels are spent covering rare extremes instead of the bulk of the distribution, which leads to a significant loss of accuracy.
+We define the transform more formally in Equation˜8.
+[…]
+2 Related Work
+Post-training quantization (PTQ) and outliers.
+PTQ schemes, such as RTN (round-to-nearest), OBQ (Frantar and Alistarh, 2022), and GPTQ (Frantar et al., 2023), are sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values can determine the quantization scale, leading to high error.
 Non-uniform bitwidths and explicit outlier storage.
 One common strategy to mitigate outliers is to use non-uniform bitwidths.
-Methods such as LLM.int8() (NEURIPS2022_c3ba4962), SpQR (dettmers2024spqr), and QUIK (ashkboos-etal-2024-quik) explicitly separate and store outliers in higher precision, while HPTQ (chen2025geometryllmquantizationgptq) uses Huffman encoding to compress the outliers and reduce their storage costs.
-These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels and additional engineering effort for efficient deployment.
+Methods such as LLM.int8() (Dettmers et al., 2022), SpQR (Dettmers et al., 2024), and QUIK (Ashkboos et al., 2024a) explicitly separate and store outliers in higher precision, while HPTQ (Chen et al., 2025) uses Huffman encoding to compress the outliers and reduce their storage costs.
+These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels.
 Transform-based outlier mitigation.
 Another line of work applies transforms to the weights and activations before quantization to reduce the impact of outliers.
-SmoothQuant (pmlr-v202-xiao23c) and AWQ (MLSYS2024_42a452cb) rescale channels to stabilize the dynamic ranges of weights and activations.
-QuIP (cheequip2023) introduces an incoherence processing step.
-QuIP# (pmlr-v235-tseng24a), QTIP (NEURIPS2024_6de2e84b), and QuaRot (NEURIPS2024_b5b93943) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (liu2025spinquant) learns rotations optimized on calibration data.
-However, their transforms are either heuristic, costly to learn, or primarily designed for weight-only quantization, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
+SmoothQuant (Xiao et al., 2023) and AWQ (Lin et al., 2024) rescale channels to stabilize the dynamic ranges of weights and activations.
+QuIP (Chee et al., 2023) introduces an incoherence processing step.
+QuIP# (Tseng et al., 2024a), QTIP (Tseng et al., 2024b), and QuaRot (Ashkboos et al., 2024b) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (Liu et al., 2025) and FlatQuant (Sun et al., 2025) learn transforms optimized on calibration data.
+However, their transforms are heuristic and costly to learn, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
 Blockwise transforms and emerging FP formats.
-The recent FP formats, such as MXFP (ocp_mx_2023) and NVFP (nvidia_blackwell_2024), have motivated blockwise transform schemes.
-shao2025blockrotationneedmxfp4 indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but egiazarian2025bridginggappromiseperformance highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions.
-In contrast, our work derives a closed-form, data-aware optimal blockwise transform (WUSH) and analyzes how such a transform interacts with both FP and INT block quantizers.
-1.3 Methodology
+The recent FP formats, such as MXFP (MX Alliance, 2023) and NVFP (NVIDIA, 2024), have motivated blockwise transform schemes.
+Shao et al.
+(2025) indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but Egiazarian et al.
+(2025) highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions.
+In contrast, this work (WUSH) derives a closed-form, data-aware optimal blockwise transform and analyzes how such a transform interacts with both FP and INT block quantizers.
+3 Methodology
 Problem setup.
@@
 For weight-activation quantization, the output activation $\bm{W}^{\top}\bm{X}$ becomes $q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)$.
-Our goal is to choose $\bm{T}_{\mathrm{W}}$, $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
+Our goal is to choose $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
 $\ell=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)-\bm{W}^{\top}\bm{X}\right\|_{\mathrm{F}}^{2}$ (2)
@@
 While the weights can be pre-quantized, the activations must be transformed and quantized dynamically at inference time.
-To reduce the computational overhead and simplify the problem, we constrain the transforms $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ to be block-diagonal, and the quantizer $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales.
-We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$ ($d$ is a factor of $d_{\mathrm{in}}$, and $d$ is a power of $2$).
-Denote
-$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right)\quad\text{and}\quad\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right)$ (3)
-with $\bm{T}_{\mathrm{W}_{\left(i\right)}},\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathrm{R}^{d\times d}$.
-We partition $\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ and $\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{X}_{\left(i\right)}\in\mathrm{R}^{d\times d_{\mathrm{batch}}}$.
-Then, we can express the output without quantization as $\bm{W}^{\top}\bm{X}=\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}$ and the output with weight-activation quantization as
-$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$ (4)
-Finally, define the blockwise output loss as
+To reduce the computational overhead of these operations, it is common (Egiazarian et al., 2025) to constrain $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}$ to be block-diagonal transforms, and $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales.
+We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$, with $d$ being a factor of $d_{\mathrm{in}}$ and a power of $2$.
+Denote the block partitions of the matrices as
+$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right),$ $\bm{T}_{\mathrm{W}_{\left(i\right)}}\in\mathbb{R}^{d\times d},$ (3)
+$\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right),$ $\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathbb{R}^{d\times d},$
+$\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right],$ $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}},$
+$\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right],$ $\bm{X}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{batch}}}.$
+Then, we can express the output without and with weight-activation quantization, respectively, as
+$\bm{W}^{\top}\bm{X}=$ $\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)},$ (4)
+$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=$ $\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$
+Finally, define the (normalized) blockwise output loss as
 $\ell_{\left(i\right)}=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right)-\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}\right\|_{\mathrm{F}}^{2}.$ (5)
-We approximate the layerwise loss as $\ell\approx\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise loss $\ell_{\left(i\right)}$ independently.
+We will approximate the layerwise loss as $\ell$ $\approx$ $\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise losses $\ell_{\left(i\right)}$ independently.
 Optimal block-diagonal transforms.
-We will prove that there exist closed-form optimal transforms for each block under reasonable assumptions and smooth approximations, which are optimal for FP and asymptotically optimal for INT quantizations.
-Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy $\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$, respectively.
-Without loss of generality (any differences will be canceled in later steps), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$.
-In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moment before the Cholesky decomposition.
-Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$.
-Let $\bm{H}\in\mathbb{R}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix where each element is $\pm d^{-\frac{1}{2}}$.
-The optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ are constructed as
-${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top}\quad\text{and}\quad{\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$ (6)
+In the following, we will prove, under reasonable assumptions, that there exist closed-form optimal transforms for each block.
+These transforms are provably optimal for FP and asymptotically optimal for INT quantization.
+Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy
+$\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=$ $\ d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top},$ (6)
+$\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=$ $\ d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top},$
 respectively.
-Note that ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}={\bm{T}_{\mathrm{hsuw}}^{-\top}}_{\left(i\right)}$  [^1] The -$\top$ superscript notation means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$.
-, which can be easily verified by calculating ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}{\bm{T}_{\mathrm{hsuw}}^{\top}}_{\left(i\right)}=\mathbf{I}$ with $\mathbf{I}$ as the identity matrix.
+Without loss of generality (Appendix Section˜A.1), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$.
+In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moments before the Cholesky decomposition.
+Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of
+$\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}.$ (7)
+Let $\bm{H}\in\bigl\{\pm d^{-\frac{1}{2}}\bigr\}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix.
+Then, we will show that the optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ can be constructed as
+${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}=$ $\ \bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top},$ (8)
+${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}=$ $\ \bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$
+respectively. Note that [^1] The -$\top$ superscript means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$.
+${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}={\bm{T}_{\mathrm{wush}}^{-\top}}_{\left(i\right)},$ (9)
+which can be easily verified by calculating ${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}{\bm{T}_{\mathrm{wush}}^{\top}}_{\left(i\right)}=\mathbf{I}$, where $\mathbf{I}$ is the identity matrix.
 Remark.
-The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Eq.
-(6), which explains why it has been empirically shown to be one of the most effective data-agnostic orthogonal transforms.
+The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Equation˜8.
+Incidentally, this explains why it has been empirically observed to be an effective data-agnostic orthogonal transform.
+1
+21:Input:
+32:weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, Hadamard $\bm{H}\in\bigl\{\pm d^{-\frac{1}{2}}\bigr\}^{d\times d}$, damping ratio $\lambda\in\mathbb{R}_{\geq 0}$, flag flagflag of whether to quantize with GPTQ
+43:Output:
+54:WUSH transform ${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\tilde{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
+65:
+[…]
+[…]
+[…]
+109:for $i\leftarrow d_{\mathrm{in}}/d$ to 11 do
+[…]
+[…]
+[…]
+[…]
+[…]
+1615:if flagflag then
+1716:  $\tilde{\bm{W}}_{\left(i\right)}\leftarrow\operatorname{GPTQ}\left(\mathrm{weight}=\bar{\bm{W}},\mathrm{Hessian}=\bm{H}\bm{S}\bm{H}^{\top}\right)$
+[…]
+1917:  $\bm{Y}\leftarrow\bm{Y}-\tilde{\bm{W}}_{\left(i\right)}^{\top}{\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\bm{X}^{\prime}[j_{1}:j_{2},1:d_{\mathrm{in}}]$ $\triangleright$ inter-block GPTQ error propagation
+2018:else
+2119:  $\tilde{\bm{W}}_{\left(i\right)}\leftarrow q\left(\bar{\bm{W}}\right)$ $\triangleright$ RTN (round-to-nearest) fallback
+2220:  $\bm{Y}\leftarrow\bm{Y}-\bar{\bm{W}}^{\top}{\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\bm{X}^{\prime}[j_{1}:j_{2},1:d_{\mathrm{in}}]$ $\triangleright$ blockwise triangular solve
+2321:end if
+2422:end for
+Algorithm 1 Compute WUSH and Pre-Quantize Weights
+GPTQ integration.
+In practice, the second-order information $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$ in Equation˜6 can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ.
+Therefore, it is a natural idea to integrate WUSH into GPTQ.
+However, GPTQ updates the weights iteratively via error propagation, while the WUSH transform is constructed from the second-order statistics of the updated weights.
+This coupling between weight updates and transform construction requires an interleaved computational schedule.
+Algorithm˜1 summarizes the procedure used to compute the blockwise [^2] The $\cdots\left[a:b,c:d\right]$ notation in algorithm selects block with row indices from $a$ to $b$ and column indices from $c$ to $d$, all inclusive.
+WUSH transforms and pre-quantized weights for a linear layer, using either GPTQ or RTN.
+The GPTQ error propagations are decomposed into intra-block and inter-block updates similar to its original paper (Frantar et al., 2023).
+For the intra-block updates, we apply the standard GPTQ subroutine to the current transformed weight block using the corresponding transformed Hessian.
+For inter-block propagation, we follow GPTQ’s usual blockwise updates, but we implement them using the equivalent formulation from Chen et al.
+(2025), which simplifies the procedure by avoiding explicit Hessian inversion.
+The model-level calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one.
+During inference, the forward pass of a linear layer is calculated as
+$\sum_{i=1}^{d_{\mathrm{in}}/d}\tilde{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\tilde{\bm{X}}_{\left(i\right)}\right)$ (10)
+with $\tilde{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ being the pre-quantized and transformed weight blocks and $\tilde{\bm{X}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{batch}}}$ being the new activation blocks instead of the calibration ones.
 4 Theoretical Derivation
-In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$. We omit the subscript $\left(i\right)$ to simplify notation.
+In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$ in Equation˜5.
+To build intuition for the proof, Figure˜1 visualizes how different transforms reshape the expected quantization error under FP and INT block quantizers, in a 2D toy setting.
+Throughout, we omit the block subscript $\left(i\right)$ to simplify notation.
+Also, the detailed derivation steps of some equations are deferred to Appendix Section˜A.3.
+Figure 1: 2D illustration of how transforms shape the one-sided quantization error $\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\|^{2}$ in Equation˜14 under FP and INT block quantizers.
+Each row corresponds to a different transform $\bm{T}$ (identity; Hadamard: 45° rotation; calibrated Hadamard: equally spreading energy into each dimension; whitening; WUS: WUSH without Hadamard; WUSH), and columns show FP and INT quantization in both original and transformed space.
+The dotted grid shows the basis induced by the transform.
+Two example points are shown in detail with their induced quantization noise support (Sections˜4.2.1 and 4.3.1) represented as shaded parallelograms, and the corresponding average (expected) error magnitude (Equation˜14) represented as the radii of circles (not necessarily proportional to the area of parallelograms).
+For each point on the plane, we visualize its expected error change (compared to that of the identity transform) calculated in the original space, with blue/red background heatmaps indicating lower/higher and dashed lines indicating the equilibrium set where the expected error is unchanged.
+No transform can reduce the error for all points on the plane, otherwise we would be able to repeatedly apply that transform until the error reaches zero.
+Thus, the key idea is to reduce the error around typical data distributions.
+The black ellipse depicts a representative data distribution contour.
+Overall, WUSH more effectively reduces the error by aligning the ellipse’s major axis to the error reduction (blue) regions.
 4.1 Problem Setup and General Proof Approach
@@
 samples from $d$-dimensional independent distributions $\bm{w}\sim\mathcal{D}_{\mathrm{w}}$ and $\bm{x}\sim\mathcal{D}_{\mathrm{x}}$, respectively.
-Then the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy $\bm{W}^{\prime}\bm{W}^{\prime\top}=\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top}$ and $\bm{X}^{\prime}\bm{X}^{\prime\top}=\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}$, respectively.
-We can reinterpret the blockwise loss as $\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}$.
+We motivate this multivariate modeling choice in Appendix Section˜A.2.
+We can reinterpret the blockwise loss in Equation˜5 as
+$\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}.$ (11)
+Similarly, we can also reinterpret  Equation˜6 so that the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy
+$\bm{W}^{\prime}\bm{W}^{\prime\top}$ $\ =\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top},$ (12)
+$\bm{X}^{\prime}\bm{X}^{\prime\top}$ $\ =\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top},$
+respectively.
 Unbiased quantization.
@@
 We further constrain $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$ such that $q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)$ is unbiased with respect to $\bm{w}^{\top}\bm{x}$.
-Then, using the first-order approximation, we can split $\ell$ into two non-negative terms:
-$\ell=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\left(\bm{T}_{\mathrm{W}}\bm{w}+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{\top}\left(\bm{T}_{\mathrm{X}}\bm{x}+\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)-\bm{w}^{\top}\bm{x}\right)^{2}$ (7)
-$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
-$\approx$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
-$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right)$
-$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$
+Then, using the first-order approximation, we can split $\ell$ in  Equation˜11 into two non-negative terms:
+$\ell=\left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$ (13)
+The detailed derivation is given in Appendix Equation˜37.
 Problem reduction.
-We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify if $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$.
+We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify that $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$.
 The two terms have similar forms and can be optimized in similar ways.
 We focus on finding the optimal $\bm{T}_{\mathrm{X}}$ for the second term.
-Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}\sim\mathcal{D}_{\mathrm{y}}$.
+Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}$ and denote its distribution as $\bm{y}\sim\mathcal{D}_{\mathrm{y}}$.
 Define the transformation matrix $\bm{T}=\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}$.
-Then, the second term becomes
+Then, the second term in Equation˜13 becomes
 $\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\left(\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}\right)^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\left(\bm{W}^{\prime-\top}\bm{y}\right)\right)\right\|^{2}$ (14)
 $=$ $\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$
-Therefore, the two-sided quantization problem is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute $\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}$.
+Therefore, the two-sided quantization problem in Equation˜11 is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute
+$\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}.$ (15)
 Transformation parameterization.
-Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$; thus, $\bm{T}$ can be parameterized as $\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}$.
+Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$ with $\bm{U},\bm{S}$ defined in Equation˜7; thus, $\bm{T}$ can be parameterized as
+$\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}.$ (16)
 Denote $\bm{S}=\operatorname{diag}\left(s_{1},\dots,s_{d}\right)$ and $\bm{S}^{\prime}=\operatorname{diag}\left(s^{\prime}_{1},\dots,s^{\prime}_{d}\right)$.
 Without loss of generality, assume $s_{1}\geq\dots\geq s_{d}>0$ and $s^{\prime}_{1}\geq\dots\geq s^{\prime}_{d}>0$.
+A useful property of $\bm{S}$ is
+$\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$ (17)
+with detailed steps in Equation˜38.
+The equality $\operatorname{tr}\left(\bm{S}^{2}\right)=d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ is attained when all $s_{i}$ are the same (all singular values are inliers).
+The equality $d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}=d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$ is attained when $\operatorname{tr}\left(\bm{S}\right)\to s_{1}$ (a singular value is an extreme outlier).
 Theorem 1 (Optimal Transform)
 The optimal configuration of $\bm{T}$ for floating-point (FP) data types is $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$.
-For integer (INT), the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
-Proof We provide detailed proofs for FP and INT in Sections˜2.3 and 2.2, respectively.
+For integer (INT) types, the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean multivariate Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
+Proof We provide detailed proofs for FP and INT in Sections˜4.2 and 4.3, respectively.
 For each data type, we apply a two-step proof strategy.
-We first provide a smooth modeling of the quantizer and treat the type casting errors as random noise.
-Secondly, we calculate the expected quantization error with respect to the parameterized transform and minimize this expectation by solving the corresponding optimization problem.
-Some useful notations and basic results are as follows:
-  • Denote the $k$-th standard basis vector as $\mathbf{e}_{k}\in\left\{0,1\right\}^{d}$, where the $k$-th element is $1$ and 0 otherwise.
-  • The second moment of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}=\bm{W}^{\prime\top}\left(\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}\right)\bm{W}^{\prime}=\bm{W}^{\prime\top}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{W}^{\prime}=\bm{U}\bm{S}^{2}\bm{U}^{\top}$.
-  • The second moment of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}=\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}=\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}$.
-  • The expected squared norm of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$.
-  • Similarly, the expected squared norm of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$.
-  • Because $\bm{S}\succ\bm{0}$, $\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\right)$.
-By Jensen’s inequality, $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$.
+We first (Sections˜4.2.1 and 4.3.1) provide a smooth modeling of the quantizer and treat the type casting errors as random noise.
+Secondly (Sections˜4.2.2 and 4.3.2), we calculate the expected quantization error, Equation˜14, with respect to the parameterized transform, Equation˜16, and minimize this expectation by solving the corresponding optimization problem.
+The optimal $\bm{T}_{\mathrm{X}}$, aka the WUSH construction $\bm{T}_{\mathrm{wush}}$ in Equation˜8, is obtained by applying Theorems˜1, 15 and 16.
 4.2 Proof for Floating-Point (FP) Types
 4.2.1 Quantization Error Modeling
-Theorem 2 (FP Quantization Error Modeling)
-For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d.
-samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
-Proof The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$.
-Assume we do not need to consider the subnormal and special number cases.
-The tuple $\left(s,e,m\right)$ represents the number $x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right)$.
-Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$, we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as $x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ with $2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ being a $a+b$ bit unsigned integer (concatenating the binary representations of $e$ and $m$).
-The minimum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ is $\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{-2^{a-1}+1}$ and the maximum is $\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1}$.
-Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type.
-Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely.
-$\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in the logarithm space and first-order expansion.
-For $x\neq 0$,
-$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}=\mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$ (9)
-$\approx$ $\ \mathrm{sgn}\left(x\right)\left(2^{2^{-b}2^{b}\log_{2}\left|x\right|}+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{2^{-b}2^{b}\log_{2}\left|x\right|}2^{-b}\ln 2\right)$
-$=$ $\ x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right).$
-For $x=0$, $\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0$.
-We model the casting error $\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
+For a number in FP types, the quantization error tends to be proportional to its absolute value. Formally, for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error as
+$\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ (18)
+where $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ is a random vector of i.i.d.
+samples from the distribution $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ with $\mathbb{E}_{\eta}\eta=0$.
+We provide more rigorous justifications for why this modeling makes sense in Appendix Section˜A.4.2.
+4.2.2 Loss Minimization
+Using the quantization error modeling in Equation˜18, the minimization objective in Equation˜14 becomes
+$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (19)
+with $\odot$ representing elementwise multiplication, and with detailed steps in Equation˜39. The lower bound of the trace term in Equation˜19 is
+$\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (20)
+and the equality can be attained by choosing $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$, with detailed steps in Equations˜40 and 41.
+4.2.3 Discussion
+Under our smooth modeling in Equation˜18, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective becomes trivially
+$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)$ (21)
+with detailed steps in Equation˜42.
+Thus, orthogonal transforms will not be helpful for reducing the quantization error.
+Comparing the trace terms in Equations˜20 and 21 using the inequality Equation˜17, the WUSH transform can reduce the error by at most $d$ times in the extreme outlier scenario.
+Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error.
+Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term
+$\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$ (22)
+as that in the case of the orthogonal $\bm{T}$ in Equation˜21, with detailed steps in Equations˜43 and 44. Figure˜1 (left two columns) also visualizes these results.
+4.3 Proof for Integer (INT) Data Types
+4.3.1 Quantization Error Modeling
+For a number in INT types, the quantization error tends to be proportional to the maximum absolute value within a quantization group.
+Formally, for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error as
+$\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ (23)
+where $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ is a random vector of i.i.d.
+samples from the distribution $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ with $\mathbb{E}_{\eta}\eta=0$.
+We provide more rigorous justifications for why this modeling makes sense in Appendix Section˜A.4.1.
+4.3.2 Loss Minimization
+Using the quantization error modeling in Equation˜23, the minimization objective in Equation˜14 becomes
+$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ (24)
+with detailed steps in Equation˜45.
+We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}$ in Equation˜24.
+$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right),$ (25)
+and the equality is attained when $\bm{R}=\mathbf{I}$, with detailed steps in Equations˜46 and 47.
+Next, we obtain near-optimal lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$ in Equation˜24 by choosing $\bm{U}^{\prime}=\bm{H}$.
 […]
-$q\left(x\right)=$ $\ \mathrm{SE}a\mathrm{M}b\left(xs^{-1}\right)\mathrm{SE}\tilde{a}\mathrm{M}\tilde{b}\left(s\right)=\left(1+\left(\ln 2\right)2^{-b}\xi\right)xs^{-1}\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)s$ (10)
-$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)x$
-with the independent random variables $\xi,\tilde{\xi}\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
-Define $\eta=\left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)-1$. The quantization error is $q\left(x\right)-x=\eta x$.
-$\mathbb{E}_{\eta}\eta=\mathbb{E}_{\xi,\tilde{\xi}}\left(\ln 2\right)2^{-b}\xi+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}+\left(\ln 2\right)^{2}2^{-b-\tilde{b}}\xi\tilde{\xi}=0.$ (11)
-2.2.2 Loss Minimization
-Using Theorem˜2, the minimization objective is
-$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (12)
+Note that the upper bound is tighter when $\mathcal{D}_{\mathrm{y}}$ is a tail-bounded (zero-mean multivariate Gaussian or Laplacian) distribution than for a general distribution.
+For the full derivations of these bounds, please refer to Eqs.
+(48) to (57).
+Taken Equations˜25 and 26 together, by setting $\bm{R}=\mathbf{I}$ and $\bm{U}^{\prime}=\bm{H}$, the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ is bounded by $\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ within a gap of $d$ such that
+$d^{-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (27)
+And for a tail-bounded $\mathcal{D}_{\mathrm{y}}$, the gap is narrowed to $d^{o\left(1\right)}$ such that
+$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (28)
+By the Cauchy-Schwarz inequality, the lower bound of the trace product term is
+$\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\geq\left(\operatorname{tr}\left(\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)^{\frac{1}{2}}\left(\bm{S}^{\prime 2}\right)^{\frac{1}{2}}\right)\right)^{2}=\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$ (29)
+and the equality is attained when $\bm{S}^{\prime}\propto\bm{S}^{\frac{1}{2}}$.
+Without loss of generality, we can choose $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ to minimize both the lower and upper bounds in Equations˜27 and 28, and the quantization loss in Equation˜24 becomes near-optimal:
+[…]
+4.3.3 Discussion
+Consider the case of $\bm{T}$ being orthogonal. For any general distribution $\mathcal{D}_{\mathrm{y}}$, the bounds are
+$\operatorname{tr}\left(\bm{S}^{2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq d\operatorname{tr}\left(\bm{S}^{2}\right)$ (31)
+with detailed steps in Equation˜58. For a tail-bounded $\mathcal{D}_{\mathrm{y}}$, the Hadamard $\bm{T}=\bm{H}$ is empirically the best orthogonal transform, so the bounds can be tightened to
+$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)}\operatorname{tr}\left(\bm{S}^{2}\right)$ (32)
+Comparing Equation˜30 with Equations˜31 and 32 using the inequality Equation˜17, an orthogonal $\bm{T}$ is suboptimal, and the error bounds of the Hadamard $\bm{T}=\bm{H}$ can be at most $d$ times larger than those of the optimal $\bm{T}$, in the extreme outlier scenario.
+Figure˜1 (right two columns) provides visualizations of different transforms.
+5 GPU Kernel Support
+As noted in Equation˜10, part of the WUSH transform has to be applied online during inference, requiring a specialized GPU kernel.
+Online blockwise Hadamard rotations are efficiently supported in MR-GPTQ’s kernels (Egiazarian et al., 2025).
+Yet, in their case, the transforms are data-independent, so it is possible to reuse a single transform across all blocks.
+By contrast, WUSH assigns a distinct transform to each block.
+This per-block specialization significantly complicates kernel reuse and fusion in existing low-level libraries.
+[…]
+[…]
+[…]
+[…]
+6 Experiments
+Setup.
+We now present experiments evaluating the layer-wise losses and the accuracy of quantized models using the WUSH transform, relative to SmoothQuant (Xiao et al., 2023), QuaRot (Ashkboos et al., 2024b), SpinQuant (Liu et al., 2025), and MR-GPTQ (Egiazarian et al., 2025) baselines.
+MR-GPTQ is the current state-of-the-art for MXFP/NVFP quantization.
+We conduct experiments on the LLama-3.2-3B-Instruct, LLama-3.1-8B-Instruct, and Qwen3-8B/14B/32B models.
+We use Platinum Benchmarks (Vendrow et al., 2025) and the LM Evaluation Harness (Gao et al., 2021) for accuracy evaluation.
+6.1 Superior Per-Layer Layer Loss
+Table 1: Layerwise RTN quantization loss in the scale of $10^{-3}$.
+Q K V O G U D
+MXFP4 I 11.1 12.0 10.7 4.35 7.10 6.56 5.47
+R 7.61 7.73 9.14 3.84 5.56 5.68 4.07
+H 7.24 7.20 8.60 3.79 5.45 5.61 3.90
+WUS 6.27 7.22 4.05 3.57 5.76 4.75 4.46
+WUSH 3.34 3.34 3.30 2.76 4.49 4.39 3.39
+NVFP4 I 4.23 4.35 4.37 2.34 3.49 3.41 2.41
+R 4.98 5.01 5.86 2.54 3.63 3.68 2.66
+H 5.60 5.62 6.70 2.58 3.71 3.79 2.78
+WUS 2.26 2.36 2.30 2.00 3.04 3.01 2.23
+WUSH 2.40 2.44 2.28 1.92 3.09 3.02 2.33
+INT4 I 170. 123. 13.2 4.55 55.9 9.83 19.3
+R 5.79 5.84 6.96 2.94 4.22 4.29 3.10
+H 5.57 5.55 6.80 2.86 4.09 4.25 3.03
+WUS 213. 142. 10.7 4.54 50.2 7.42 13.1
+WUSH 2.39 2.43 2.54 2.10 3.43 3.43 2.55
+We first report the layerwise weight-activation RTN quantization loss ($L_{2}$ loss normalized by the number of elements) for each linear layer (attention projection layers Q, K, V, O and MLP projection layers Gate (G), Up (U), Down (D)) in the 18th transformer block of Qwen3-8B, using 32 calibration samples of sequence length 2048 from the FineWeb-Edu dataset (Penedo et al., 2024).
+We found the results to be consistent across blocks and datasets.
+We compare the WUSH transform with the identity (I), random rotation (R) averaged over 10 runs, Hadamard (H), and WUSH without Hadamard (WUS).
+Table˜1 summarizes the losses for MXFP4, NVFP4, and INT4 formats, where INT4 denotes 4-bit integers with Gaussian MSE clipping and BF16 scales of group size 32.
+Across formats, WUSH substantially reduces the loss: WUSH consistently yields the smallest loss for MXFP4 and INT4, while WUSH and WUS are almost equally optimal for NVFP4.
+The errors for NVFP4 tend to be lower due to its group size of 16.
+These trends confirm our theory, suggesting that the assumptions (Section˜4.1) and modeling (Sections˜4.2.1 and 4.3.1) capture the dominant sources of error.
+At the same time, MXFP4 behaves as a hybrid between ideal FP and INT quantization: the effective mantissa step changes uniformly, imparting INT-like behavior, especially in the subnormal regime and small exponent ranges.
+This explains why Hadamard transforms provide measurable gains for MXFP4, despite the ideal FP theory (Section˜4.2.3) predicting no benefit for purely orthogonal transforms.
+For NVFP4, the Hadamard alone is harmful due to the top-element preservation effect of this format (Egiazarian et al., 2025).
+Remarkably, other components in WUSH overcome this effect and produce smaller losses than the identity transform.
+6.2 End-to-end LLM Accuracy Benchmarks
+Table 2: Llama-3.1-8B-Instruct W4A4 accuracy results on the LM Evaluation Harness under different methods.
+Format Method MMLU-CoT GSM8K HellaSwag WinoGrande Average Recovery
+BF16 - 72.76 85.06 80.01 77.90 78.93 100.0
+NVFP4 RTN-I 68.26 78.39 78.15 74.11 74.73 94.67
+RTN-H 67.41 78.01 77.31 73.48 74.05 93.82
+SmoothQuant 68.90 79.50 79.50 74.70 75.70 95.90
+QuaRot 66.50 77.40 77.25 75.14 74.10 93.80
+SpinQuant 66.50 76.10 76.96 75.32 73.70 93.40
+GPTQ-I 68.85 82.60 78.26 74.51 75.72 95.92
+GPTQ-H (MR-GPTQ) 69.12 80.80 78.17 75.24 75.84 96.08
+RTN-WUSH 68.83 78.57 78.22 75.47 75.28 95.37
+GPTQ-WUSH 69.69 80.11 78.52 76.09 76.10 96.40
+MXFP4 RTN-I 62.21 67.85 73.99 73.24 69.32 87.83
+RTN-H 62.38 72.48 75.29 71.67 70.45 89.26
+SmoothQuant 63.93 68.54 75.10 73.56 70.30 89.06
+QuaRot 49.86 56.94 73.50 71.43 62.90 79.70
+SpinQuant 61.80 68.16 74.87 72.93 69.40 88.00
+GPTQ-I 63.49 68.46 76.01 74.51 70.62 89.47
+GPTQ-H (MR-GPTQ) 67.19 75.70 76.91 74.80 73.65 93.31
+RTN-WUSH 66.85 75.16 77.28 73.56 73.21 92.75
+GPTQ-WUSH 67.79 77.41 77.44 74.78 74.35 94.20
+Figure 3: Comparison of different transforms on Qwen3-8B.
+Table˜2 compares WUSH relative to prior works for Llama-3.1-8B-Instruct on LM Evaluation Harness metrics, in the same setup as Egiazarian et al.
+(2025).
+Overall, WUSH consistently yields the best results, by a large margin for MXFP4.
+On Qwen3-8B measured on Platinum Benchmarks (Figure˜3), WUSH improves MXFP4 by up to +1.3 (RTN) and +0.7 (GPTQ) average points, while NVFP4 improvements are smaller and often within run-to-run variability.
+Results (Tables˜4, 8, 3, 9, 6 and 7) for additional models are in Appendix Appendix˜B.
+Table˜4 shows the LM Evaluation Harness results for different models, e.g., for Qwen3-14B WUSH reduces the NVFP4-MXFP4 average gap to 0.36 points, and above 98% recovery.
+Figures˜8, 9, 6 and 7 also reflect similar trends.
+We conclude that WUSH improves the state-of-the-art for weight-activation quantization.
+7 Future Work
+It will be useful to explore structurally simplified variants of WUSH that trade optimality for a reduction in the cost of applying the transforms at inference time.
+One promising avenue is to approximate the data-aware component with a diagonal matrix.
+8 Conclusion
+We have studied weight-activation quantization for LLMs in the presence of heavy-tailed outliers and have shown that this problem admits closed-form linear transforms for both FP and INT block quantizers.
+Our construction, WUSH, combines a fixed Hadamard backbone with a data-aware component derived from second-order statistics, yielding a non-orthogonal blockwise transform that is provably near-optimal and remains amenable to efficient GPU kernel implementations.
+These results clarify the empirical success of Hadamard rotations and indicate that principled, data-aware block transforms can yield significant improvements, e.g., making the MXFP format competitive with NVFP.
+Acknowledgment
+We would like to thank Tijmen Blankevoort for useful discussions and for suggesting the name WUSH, which is a clear improvement over our previous name (HSUW).
+Appendix A Detailed Theoretical Results
+A.1 Second Moment Decompositions
+Lemma 2 (Invariance of the Decomposition Choices)
+For any full rank $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}$ that satisfy Equation˜6, the matrix products $\bm{W}_{\left(i\right)}^{\prime}\bm{U}_{\left(i\right)}$ and $\bm{X}_{\left(i\right)}^{\prime}\bm{V}_{\left(i\right)}$ in Equation˜8 are invariant up to sign and permutation changes.
+Proof We omit the subscript $\left(i\right)$ and superscript to simplify notation.
+Let $\hat{\bm{W}},\hat{\bm{X}}\in\mathbb{R}^{d\times d}$ such that $\hat{\bm{W}}\hat{\bm{W}}^{\top}=\bm{W}\bm{W}^{\top}$ and $\hat{\bm{X}}\hat{\bm{X}}^{\top}=\bm{X}\bm{X}^{\top}$.
+Define $\bm{Q}_{\mathrm{W}}=\bm{W}^{-1}\hat{\bm{W}}$ and $\bm{Q}_{\mathrm{X}}=\bm{X}^{-1}\hat{\bm{X}}$.
+Then,
+$\bm{Q}_{\mathrm{W}}\bm{Q}_{\mathrm{W}}^{\top}=\bm{W}^{-1}\hat{\bm{W}}\hat{\bm{W}}^{\top}\bm{W}^{-\top}=\bm{W}^{-1}(\bm{W}\bm{W}^{\top})\bm{W}^{-\top}=\mathbf{I},$ (33)
+and similarly $\bm{Q}_{\mathrm{X}}\bm{Q}_{\mathrm{X}}^{\top}=\mathbf{I}$.
+Hence $\bm{Q}_{\mathrm{W}},\bm{Q}_{\mathrm{X}}$ are orthogonal and $\hat{\bm{W}}=\bm{W}\bm{Q}_{\mathrm{W}}$, $\hat{\bm{X}}=\bm{X}\bm{Q}_{\mathrm{X}}$.
+By Equation˜7, we have $\bm{U},\bm{S},\bm{V}\in\mathbb{R}^{d\times d}$ as the SVD of $\bm{W}^{\top}\bm{X}=\bm{U}\bm{S}\bm{V}^{\top}$.
+Therefore,
+$\hat{\bm{W}}^{\top}\hat{\bm{X}}=\bm{Q}_{\mathrm{W}}^{\top}\bm{W}^{\top}\bm{X}\bm{Q}_{\mathrm{X}}=\bm{Q}_{\mathrm{W}}^{\top}\left(\bm{U}\bm{S}\bm{V}^{\top}\right)\bm{Q}_{\mathrm{X}}=\left(\bm{Q}_{\mathrm{W}}^{\top}\bm{U}\right)\bm{S}\left(\bm{Q}_{\mathrm{X}}^{\top}\bm{V}\right)^{\top}.$ (34)
+Thus, we may take $\hat{\bm{U}}=\bm{Q}_{\mathrm{W}}^{\top}\bm{U}$, $\hat{\bm{S}}=\bm{S}$, $\hat{\bm{V}}=\bm{Q}_{X}^{\top}\bm{V}$ (up to the sign and permutation conventions) as the SVD of $\hat{\bm{W}}^{\top}\hat{\bm{X}}=\hat{\bm{U}}\hat{\bm{S}}\hat{\bm{V}}^{\top}$.
+With this,
+$\hat{\bm{W}}\hat{\bm{U}}$= $\ \bm{W}\bm{Q}_{\mathrm{W}}\bm{Q}_{\mathrm{W}}^{\top}\bm{U}=\bm{W}\bm{U},$ (35)
+$\hat{\bm{X}}\hat{\bm{V}}$= $\ \bm{X}\bm{Q}_{\mathrm{X}}\bm{Q}_{\mathrm{X}}^{\top}\bm{V}=\bm{X}\bm{V}.$
+A.2 Multivariate vs. Univariate Probabilistic Modeling
+A common intuition in prior transform-based quantization methods is largely univariate: apply an orthogonal mixing (Hadamard or learned rotation) so that the distribution of coordinate-wise values becomes more “Gaussian-like,” thereby reducing the impact of heavy-tailed outliers.
+In contrast, our probabilistic modeling is explicitly multivariate: we model a quantization group as a joint random vector and reason about its second-order structure, enabling analysis in the eigenbasis and construction of transforms that depend on the spectrum (eigenvalues / singular values).
+Figure 4: Rotation can equalize marginals without alleviating multivariate outliers.
+Top: an anisotropic 2D distribution with unequal coordinate marginals.
+Bottom: after an orthogonal rotation, the two marginals become nearly identical, and their average appears more Gaussian-like.
+Nevertheless, the joint distribution retains the same eigenvalues (spectrum) under rotation, so extreme points remain extreme in the multivariate space; they are merely “hidden” from any single marginal view.
+This motivates multivariate, spectrum-aware transforms that act on the joint geometry rather than relying on marginal “Gaussianization.”
+Figure˜4 illustrates the key limitation of purely orthogonal “Gaussianization.” In the top row (original distributions), the joint density is anisotropic: the two coordinate-wise marginals differ in spread, and their average can appear heavy-tailed or non-Gaussian depending on the axis alignment.
+In the bottom row, after the Hadamard ($45^{\circ}$ rotation) transform, the coordinate marginals become nearly identical, and the average marginal appears more Gaussian-like.
+However, this does not imply that outliers have been removed in the underlying multivariate geometry: orthogonal transforms preserve global energy and, crucially, preserve the eigenvalues of the covariance (they only rotate eigenvectors).
+Consequently, points that are extreme in the principal directions remain extreme in the joint space, even if they are less visually apparent in any single marginal.
+This distinction is central to group quantization.
+Quantization error is driven by joint properties (e.g., how mass concentrates along dominant directions and how extreme points interact with group-wise scaling), which cannot be certified by marginal statistics alone.
+Our multivariate formulation therefore targets the spectrum directly: by exposing and using the eigen-/singular-value structure of the group statistics, we can design transforms that genuinely reduce the effective anisotropy responsible for outlier-dominated scaling, rather than merely redistributing it across coordinates via rotations.
+A.3 Equation Derivations
+Standard basis vector notation.
+Denote the $k$-th standard basis vector as $\mathbf{e}_{k}\in\left\{0,1\right\}^{d}$, where the $k$-th element is $1$ and 0 otherwise.
+Basic second-order expectations.
+$\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}=\bm{W}^{\prime\top}\left(\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}\right)\bm{W}^{\prime}=\bm{W}^{\prime\top}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{W}^{\prime}$ (Equation˜12) (36)
+$\quad\quad\quad\quad=\bm{U}\bm{S}\bm{V}^{\top}\bm{V}\bm{S}\bm{U}^{\top}=\bm{U}\bm{S}^{2}\bm{U}^{\top},$ (Equation˜7)
+$\mathbb{E}_{\bm{y}}\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}=\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}$
+$\quad\quad\quad\quad=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}\bm{U}\bm{S}^{2}\bm{U}^{\top}\bm{U}\bm{S}^{-1}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}=\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top},$ (Equation˜16)
+$\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right),$
+$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}\right)=\operatorname{tr}\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)=\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$
+Equation˜13.
+$\ell=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\left(\bm{T}_{\mathrm{W}}\bm{w}+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{\top}\left(\bm{T}_{\mathrm{X}}\bm{x}+\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)-\bm{w}^{\top}\bm{x}\right)^{2}$ (37)
+$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}$
+$\quad\Bigl(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)+\underbrace{\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}_{\text{\makebox[0.0pt]{$\approx 0$ \quad(first order approx.)}}}+\underbrace{\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\bm{T}_{\mathrm{X}}\bm{x}-\bm{w}^{\top}\bm{x}}_{\text{\makebox[0.0pt]{$=0$ \quad($\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$)}}}\Bigr)^{2}$
+$\approx$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
+$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}$
+$\quad\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{2}+\left(\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}+\bigl(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\underbrace{\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}_{\text{\makebox[0.0pt]{$\mathbb{E}\left(\cdot\right)=0$}}}\bigr)\bigl(\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\underbrace{\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}_{\text{\makebox[0.0pt]{$\mathbb{E}\left(\cdot\right)=0$}}}\bigr)$
+$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{2}+\left(\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
+$=$ $\ \Bigl(\mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\bm{T}_{\mathrm{X}}\underbrace{\bm{x}\bm{x}^{\top}}_{\text{\makebox[0.0pt]{$\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}=\bm{X}^{\prime}\bm{X}^{\prime\top}$}}}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\Bigr)$
+$\quad+\Bigl(\mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)^{\top}\bm{T}_{\mathrm{W}}\underbrace{\bm{w}\bm{w}^{\top}}_{\text{\makebox[0.0pt]{$\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top}=\bm{W}^{\prime}\bm{W}^{\prime\top}$}}}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\Bigr)$
+$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\bm{T}_{\mathrm{X}}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)$
+$\quad+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)^{\top}\bm{T}_{\mathrm{W}}\bm{W}^{\prime}\bm{W}^{\prime\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)$
+$=$ $\ \Bigl(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\bigl\|\bm{X}^{\prime\top}\underbrace{\bm{T}_{\mathrm{X}}^{\top}}_{\text{\makebox[0.0pt]{$=\bm{T}_{\mathrm{W}}^{-1}$}}}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\bigr\|^{2}\Bigr)+\Bigl(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\bigl\|\bm{W}^{\prime\top}\underbrace{\bm{T}_{\mathrm{W}}^{\top}}_{\text{\makebox[0.0pt]{$=\bm{T}_{\mathrm{X}}^{-1}$}}}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\bigr\|^{2}\Bigr)$
+$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$
+Equation˜17.
+$\left\{\begin{array}[]{lr}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\right)&\ \text{($\bm{S}\succ\bm{0}$)}\\ \operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}&\ \text{(Jensen's inequality)}\end{array}\right$.
+(38)
+$\Rightarrow\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right).$
+Equation˜19.
+$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ (39)
+$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (Equation˜18)
+$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}$
 $=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\operatorname{tr}\left(\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right)$
@@
 $=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\right)\bm{T}^{-\top}\right)$
-$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
-where $\odot$ represents elementwise multiplication. The lower bound of the trace term in the loss is
+$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\operatorname{diag}\left(\bm{\eta}\right)\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\operatorname{diag}\left(\bm{\eta}\right)\right)\bm{T}^{-\top}\right)$ (Equation˜36)
+$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\mathbb{E}_{\eta}\eta^{2}\right)\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
+$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right).$
+Equation˜20.
 $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (40)
-$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)\bm{T}^{-\top}\right)$
-$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
-$=$ $\ \sum_{k=1}^{d}\left\|{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
-$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
+$=$ $\ \operatorname{tr}\biggl(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)$ $\!\!\!\!\!\bm{T}^{-\top}\biggr)$
+$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}\bm{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
+$=$ $\ \sum_{k=1}^{d}\left\|\bm{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
+$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (Equation˜16)
 $=$ $\ \sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (rotation-invariance of $L_{2}$ norm)
@@
 $=$ $\ d\left(d^{-1}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}\right)$
-$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen)
+$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen’s inequality)
 $=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\right)\right)^{2}$
@@
 The equality can be attained by choosing $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$ such that
-$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (14)
-$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
+$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (41)
+$=$ $\ \sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (Equation˜40)
+$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}$ (plug in)
+$=$ $\ \sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}$
+$=$ $\ \sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
+$=$ $\ d\left(d^{-1}\sum_{j=1}^{d}s_{j}\right)^{2}$
 $=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$
-2.2.3 Discussion
-Under our smooth modeling in Theorem˜2, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective
-$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$ (15)
-$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right),$
-so the orthogonal transforms will not be helpful for reducing the quantization error.
-Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error by at most $d$ times, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario.
-Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term $\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$, which is the same as the case of orthogonal $\bm{T}$.
-2.3 Proof for Integer (INT) Data Types
-2.3.1 Quantization Error Modeling
-We model the quantization error using pseudo noise similar to DiffQ (defossez2022differentiable).
-Theorem 3 (INT Quantization Error Modeling)
-For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d.
-samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
-Proof Define the scalar quantizer of $b$-bit symmetric integer as $q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s$, where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$.
-To simplify the analysis, we model the quantizer as $q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
-For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$.
-Define $\eta=2^{1-b}\xi$, then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$.
-The quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d.
-samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$.
-2.3.2 Loss Minimization
-Using Theorem˜3, the minimization objective is
-$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (16)
-$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
+Equation˜21.
+$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ (42)
+$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (Equation˜18)
+$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\bm{y}^{\top}\bm{T}^{\top}\mathrm{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\mathrm{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}$
+$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\bm{y}^{\top}\bm{T}^{\top}\mathrm{diag}\left(\bm{\eta}\right)^{2}\bm{T}\bm{y}$ (orthogonality of $\bm{T}$)
+$=$ $\ \mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$
+$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\bm{T}\bm{y}$
+$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{y}$ (orthogonality of $\bm{T}$)
+$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}$
+[…]
+In the case of $\bm{U}^{\prime}=\mathbf{I}$:
+$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (43)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\left(\bm{S}^{\prime 2}\odot\mathbf{I}\right)\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (plug in)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{S}^{\prime 2}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\right).$
+In the case of $\bm{S}^{\prime}=\mathbf{I}$:
+$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (44)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (plug in)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\right).$
+Equation˜24.
+$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ (45)
+$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (Equation˜23)
+$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\bm{\eta}^{\top}\bm{T}^{-\top}\bm{T}^{-1}\bm{\eta}$
+$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\operatorname{tr}\left(\bm{\eta}^{\top}\bm{T}^{-\top}\bm{T}^{-1}\bm{\eta}\right)$
+$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\operatorname{tr}\left(\bm{T}^{-1}\bm{\eta}\bm{\eta}^{\top}\bm{T}^{-\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
+$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
 $=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}.$
-We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$.
-Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality, we obtain:
+Equation˜25.
+$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)$ (46)
+$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16
+$=$ $\ \operatorname{tr}\left(\bm{S}\bm{U}^{\top}\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\right)$
+$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$
+Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality,
 $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)\geq\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right).$ (47)
-Equality is attained when $\bm{R}=\mathbf{I}$.
-Next, we obtain the lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$.
-Express the expectations as $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$.
+The equality can be attained by choosing $\bm{R}=\mathbf{I}$.
+Express the expectations as
+$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) (48)
+and
+$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}.$ (49)
 By the relationships among the mean, maximum, and summation values, we have:
 $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq d\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},$ (50)
-where $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ and $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$.
+where, by Equation˜36,
+$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ (51)
+and
+$\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}.$ (52)
 Lemma 3 (Maximum Inequalities for Tail-Bounded Distributions)
@@
 Both inequalities hold without the need for independence.
-Proof We provide proofs for Gaussian and Laplacian distributions in Sections˜A.1 and A.2, respectively.
+Proof We provide proofs for the Gaussian and Laplacian distributions in Sections˜A.5.1 and A.5.2, respectively.
 Using Lemma 3, for a tail-bounded $\mathcal{D}_{\mathrm{y}}$ that is a zero-mean multivariate Gaussian or Laplacian, we can further tighten the upper bound by
 $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\min\left\{d^{o\left(1\right)}\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\right\}.$ (55)
-When $\bm{U}^{\prime}=\bm{H}$, the marginal second moment becomes $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)=d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$, which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$.
-Taken together, by setting $\bm{R}=\mathbf{I}$ and $\bm{U}^{\prime}=\bm{H}$, the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ is bounded by $\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ within a gap of $d$ such that
-$d^{-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (22)
-And for a tail-bounded $\mathcal{D}_{\mathrm{y}}$, the gap is narrowed to $d^{o\left(1\right)}$ such that
-$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (23)
-By the Cauchy-Schwarz inequality,
-$\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\geq\left(\operatorname{tr}\left(\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)^{\frac{1}{2}}\left(\bm{S}^{\prime 2}\right)^{\frac{1}{2}}\right)\right)^{2}=\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (24)
-and the equality is attained when $\bm{S}^{\prime}\propto\bm{S}^{\frac{1}{2}}$.
-Without loss of generality, we can choose $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ to minimize both the lower and upper bounds of the quantization loss.
-2.3.3 Discussion
-Consider the case of $\bm{T}$ being orthogonal.
-$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$.
-For any distribution $\mathcal{D}_{\mathrm{y}}$, the bounds are $\operatorname{tr}\left(\bm{S}^{2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq d\operatorname{tr}\left(\bm{S}^{2}\right)$.
-For the Hadamard $\bm{T}=\bm{H}$ and a tail-bounded $\mathcal{D}_{\mathrm{y}}$, $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)}\operatorname{tr}\left(\bm{S}^{2}\right)$ with high probability.
-Because $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$, an orthogonal $\bm{T}$ is suboptimal, and the error bounds of an orthogonal $\bm{T}$ can be at most $d$ times larger than those of the optimal $\bm{T}$, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario.
-3 Experiments
-3.1 Algorithm
-1
-Input: weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, damping ratio $\lambda\in\mathbb{R}$
-Output: WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\hat{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
-2
+When $\bm{U}^{\prime}=\bm{H}$, using Equations˜36 and 52, the marginal second moment in Equation˜55 becomes
+$\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=$ $\ \left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ (56)
+$=$ $\ d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2},$
+which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$.
+Finally, combining Equations˜50, 55 and 56,
 […]
-3
+For orthogonal $\bm{T}$, we have $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}$.
+Using Equation˜36, $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$.
+By Equation˜50,
+$\ \operatorname{tr}\left(\bm{S}^{2}\right)=d^{-1}\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$ (58)
+$\leq$ $\ \left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=d\operatorname{tr}\left(\bm{S}^{2}\right).$
+A.4 Quantization Error Modeling Justifications
+A.4.1 Integer (INT) Types
+We model the quantization error using pseudo noise similar to DiffQ (Défossez et al., 2022).
+Define the scalar quantizer of $b$-bit symmetric integer as
+$q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s,$ (59)
+where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$.
+To simplify the analysis, we model the quantizer as
+$q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ (60)
+with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$. Then, the quantization error can be expressed as
+$q\left(x\right)-x=s\xi.$ (61)
+For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$. Define
+$\eta=2^{1-b}\xi,$ (62)
+then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$. The quantization error of vector $\bm{\alpha}$ is
+$\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ (63)
+with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and
+$\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$ (64)
+A.4.2 Floating-Point (FP) Types
+The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$.
+Assume we do not need to consider the subnormal and special number cases.
+The tuple $\left(s,e,m\right)$ represents the number
+$x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right).$ (65)
+Figure 5: The two curves $y=1+x$ and $y=2^{x}$ are close for $0\leq x\leq 1$.
+Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$ (visualized in Figure˜5), we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as
+$x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ (66)
+with
+$2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ (67)
+being an $\left(a+b\right)$-bit unsigned integer (concatenating the binary representations of $e$ and $m$). The minimum and maximum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ are
+$\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=$ $2^{-2^{a-1}+1},$ (68)
+$\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=$ $2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1},$
+respectively.
+Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type.
+Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely.
+$\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in logarithmic space and first-order expansion.
+For $x\neq 0$,
+$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}$ (69)
+$=$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$
+$\approx$ $\ \mathrm{sgn}\left(x\right)\left(2^{2^{-b}2^{b}\log_{2}\left|x\right|}+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{2^{-b}2^{b}\log_{2}\left|x\right|}2^{-b}\ln 2\right)$
+$=$ $\ x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right).$
+For $x=0$,
+$\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0.$ (70)
+We model the casting error as
+$\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ (71)
+where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and, similar to the INT types in Equation˜60, the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
 […]
-5for $i\leftarrow 1\ \mathrm{to}\ d_{\mathrm{in}}/d\ \mathrm{(in\ parallel)}$ do
-6
-7 $j_{1},j_{2}\leftarrow\left(i-1\right)d+1,id$
-[…]
-8
-[…]
-9
- $\bm{W}^{\prime}\leftarrow\operatorname{Cholesky}\left(\bm{M}_{\mathrm{W}}\right)$ // $\bm{W}^{\prime}\bm{W}^{\prime\top}=\bm{M}_{\mathrm{W}}$
-10
- $\bm{U},\bm{\Lambda}\leftarrow\operatorname{Eigendecomposition}\left(\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}\right)$ // $\bm{U}\bm{\Lambda}\bm{U}^{\top}=\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}$
-11
- ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\leftarrow\bm{H}\bm{\Lambda}^{-\frac{1}{4}}\bm{U}^{\top}\bm{W}^{\prime\top}$ // WUSH transform
-12
- $\hat{\bm{W}}_{\left(i\right)}=q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}^{-\top}\bm{W}[j_{1}:j_{2},1:d_{\mathrm{out}}]\right)$ // pre-quantize weight block
-13
-14 end for
-15
-Algorithm 1 Compute WUSH Transforms and Pre-Quantize Weights
-Algorithm˜1 summarizes the procedure used to compute the blockwise transforms and pre-quantized weights for each linear layer.
-In practice, the required second-order information can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ.
-As a result, the calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one.
-During inference, the forward pass is calculated as
-$\sum_{i=1}^{d_{\mathrm{in}}/d}\hat{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\bm{X}_{\left(i\right)}\right).$ (25)
-with the symbol $\bm{X}$ being overloaded as the new activations instead of the calibration ones.
-3.2 Layer Loss
-Table 1: Weight-activation RTN quantization losses with identity (I), random rotation (R), Hadamard (H), WUSH without Hadamard (WUS), and WUSH transforms.
-Format Transform Q K V O Gate Up Down
-MXFP4 I .01105 .01201 .01068 .00435 .00710 .00656 .00547
-R .00761 .00773 .00914 .00384 .00556 .00568 .00407
-H .00724 .00720 .00860 .00379 .00545 .00561 .00390
-WUS .00627 .00722 .00405 .00357 .00576 .00475 .00446
-WUSH .00334 .00334 .00330 .00276 .00449 .00439 .00339
-NVFP4 I .00423 .00435 .00437 .00234 .00349 .00341 .00241
-R .00498 .00501 .00586 .00254 .00363 .00368 .00266
-H .00560 .00562 .00670 .00258 .00371 .00379 .00278
-WUS .00226 .00236 .00230 .00200 .00304 .00301 .00223
-WUSH .00240 .00244 .00228 .00192 .00309 .00302 .00233
-INT4 I .17004 .12277 .01315 .00455 .05585 .00983 .01934
-R .00579 .00584 .00696 .00294 .00422 .00429 .00310
-H .00557 .00555 .00680 .00286 .00409 .00425 .00303
-WUS .21268 .14233 .01067 .00454 .05021 .00742 .01308
-WUSH .00239 .00243 .00254 .00210 .00343 .00343 .00255
-We report the layerwise weight-activation RTN quantization loss ($L_{2}$ loss normalized by the number of elements) for each linear layer in the 18th transformer block of Qwen3-8B, using 32 calibration samples of sequence length 2048 from the FineWeb-Edu (NEURIPS2024_370df50c) dataset.
-We compare the WUSH transform with the identity (I), random rotation (R) averaged over 10 runs, Hadamard (H), and WUSH without Hadamard (WUS).
-Table˜1 summarizes the losses for MXFP4, NVFP4, and INT4 formats, where INT4 denotes 4-bit integers with Gaussian MSE clipping and BF16 scales of group size 32.
-Across formats, WUSH substantially reduces the loss: WUSH consistently yields the smallest loss for MXFP4 and INT4, while WUSH and WUS are almost equally optimal for NVFP4.
-These empirical trends are broadly consistent with our theoretical analysis, suggesting that the approximation assumptions (Section˜2.1) and the smooth error models (Sections˜2.2.1 and 2.3.1) capture the dominant sources of quantization loss in practice.
-At the same time, MXFP4 behaves as a hybrid between ideal FP and INT quantization: the effective mantissa step changes uniformly, imparting INT-like behavior, especially in the subnormal regime and small exponent ranges.
-This helps explain why Hadamard transforms still provide measurable gains for the MXFP4 format, despite the ideal FP theory (Section˜2.2.3) predicting no benefit for purely orthogonal transforms.
-3.3 LLM Benchmarks
-We now present preliminary experiments evaluating the accuracy of models using the WUSH transform, relative to the identity (standard RTN) and Hadamard.
-We conduct our experiment on the LLama-3.2-3B-Instruct, LLama3.1-8B-Instruct, Qwen3-8B, and Qwen3-14B models.
-We use Platinumbench (vendrow2025large) and the LM Eval Harness (gao2021framework) for evaluation.
-The full results for LM Eval Harness tasks are shown in Table˜2.
-Here, we observe that WUSH achieves the highest average recovery score among all the methods tested, often by large margins.
-Remarkably, on large models such as Qwen3-14B, it helps narrow the accuracy gap between MXFP4 and NVFP4 to nearly 0.5%.
-We plot the results for Platinumbench in Figures˜3, 4, 1 and 2. Again, we observe that WUSH provides significant accuracy improvements, especially in the case of MXFP.
-Table 2: Accuracy results on the LM Eval Harness for round-to-nearest (RTN) quantization of weights and activations with identity (I), Hadamard (H), and WUSH transforms.
-Model Format Transform MMLU GSM8K HellaSwag WinoGrande Avg. Recovery%
-Llama-3.2-3B-Instruct BF16 - 64.43 78.01 73.42 70.09 71.49
-NVFP4 I 61.02 71.04 70.90 66.77 67.43 94.33
-H 59.91 64.82 69.77 65.59 65.02 90.95
-WUSH 61.25 71.78 71.08 67.18 67.82 94.87
-MXFP4 I 56.81 60.80 67.30 64.56 62.37 87.25
-H 55.99 61.11 68.36 64.09 62.39 87.27
-WUSH 58.45 68.60 69.99 66.54 65.89 92.17
-Llama-3.1-8B-Instruct BF16 - 72.76 85.06 80.01 77.90 78.93
-NVFP4 I 68.64 80.29 77.97 74.43 75.33 95.44
-H 66.96 77.41 77.29 74.35 74.00 93.75
-WUSH 68.83 78.57 78.22 75.47 75.28 95.37
-MXFP4 I 62.21 69.52 73.76 72.61 69.53 88.08
-H 63.03 75.44 75.51 71.90 71.47 90.54
-WUSH 66.85 75.16 77.28 73.56 73.21 92.75
-Qwen3-8B BF16 - 72.98 90.90 75.52 70.56 77.49
-NVFP4 I 70.47 88.40 74.44 69.53 75.71 97.70
-H 70.19 86.35 73.02 68.11 74.42 96.03
-WUSH 70.86 89.33 74.82 69.79 76.20 98.33
-MXFP4 I 67.69 84.23 71.24 67.40 72.64 93.74
-H 67.53 87.04 71.48 67.64 73.42 94.75
-WUSH 70.29 87.95 74.06 69.50 75.45 97.36
-Qwen3-14B BF16 - 77.18 91.96 79.84 74.27 80.81
-NVFP4 I 75.20 90.22 77.38 73.32 79.03 97.80
-H 74.98 92.04 77.76 72.38 79.29 98.12
-WUSH 75.43 91.78 78.82 73.18 79.80 98.75
-MXFP4 I 72.92 90.22 76.68 71.51 77.83 96.31
-H 73.13 87.41 76.29 71.67 77.13 95.44
-WUSH 74.56 91.87 78.18 73.15 79.44 98.30
-Figure 1: Comparison of different transforms on Qwen3-8B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks.
-The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
-Figure 2: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks.
-The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
-Figure 3: Comparison of different transforms on Llama3.2-3B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks.
-The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
-Figure 4: Comparison of different transforms on Llama3.1-8B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks.
-The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
-4 Future Work
-A natural next step is to integrate WUSH with advanced post-training quantization schemes such as GPTQ.
-In this setting, the weights are updated iteratively through error propagation, while WUSH is constructed from the second-order statistics of the updated weights.
-Designing an algorithm that jointly updates the weights and the associated blockwise transforms is non-trivial and will require adapting both the optimization schedule and the calibration pipeline.
-From a systems perspective, efficient kernel support for WUSH is another important direction.
-Unlike blockwise Hadamard rotations, which reuse a single transform across all blocks, WUSH generally assigns a distinct transform to each block.
-This per-block specialization complicates kernel reuse and fusion in existing low-level libraries.
-Finally, it will be useful to explore structurally simplified variants of WUSH that trade optimality for lower computational overhead.
-One promising avenue is to approximate the data-aware component with a diagonal matrix, thereby reducing the cost of applying the transforms at inference time.
-5 Conclusion
-We have studied blockwise weight-activation quantization for LLMs in the presence of heavy-tailed outliers and have shown that, under mild assumptions, this problem admits closed-form linear transforms for both FP and INT block quantizers.
-Our construction, WUSH, combines a fixed Hadamard backbone with a data-aware component derived from second-order statistics, yielding a non-orthogonal blockwise transform that is provably optimal and remains amenable to efficient implementation.
-These results clarify the empirical success of Hadamard rotations and indicate that principled, data-aware block transforms can significantly narrow the gap between low-bit and full-precision models.
-Acknowledgment
-We would like to thank Tijmen Blankevoort for useful discussions and for suggesting the name WUSH, which is a clear improvement over our previous name (HSUW).
-Appendix A Maximum Inequalities for Tail-Bounded Distributions
+$q\left(x\right)=$ $\ \mathrm{SE}a\mathrm{M}b\left(xs^{-1}\right)\mathrm{SE}\tilde{a}\mathrm{M}\tilde{b}\left(s\right)$ (72)
+$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)xs^{-1}\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)s$
+$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)x$
+with the independent random variables $\xi,\tilde{\xi}\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
+Define
+$\eta=\left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)-1.$ (73)
+Then, the quantization error can be expressed as
+$q\left(x\right)-x=\eta x.$ (74)
+And the expected quantization error is 0 because
+$\mathbb{E}_{\eta}\eta=\mathbb{E}_{\xi,\tilde{\xi}}\left(\ln 2\right)2^{-b}\xi+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}+\left(\ln 2\right)^{2}2^{-b-\tilde{b}}\xi\tilde{\xi}=0.$ (75)
+A.5 Maximum Inequalities for Tail-Bounded Distributions
 Lemma 4 (Expectation from the Survival Function)
-Let $X\geq 0$ be a random variable with the density function $f$ and the survival function $S(t)=\mathbb{P}\left(X\geq t\right)$ with $t\geq 0$.
-Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define $G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u$.
-Assume that $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$.
-Then $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
-Proof Since $X$ has density $f$ and takes values in $0,\infty)$, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t$.
-The survival function satisfies $S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f(u)\mathrm{d}u$, so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$.
-Hence, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t$.
+Let $X\geq 0$ be a random variable with the density function $f$ and the survival function
+$S(t)=\mathbb{P}\left(X\geq t\right)$ (76)
+with $t\geq 0$. Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define
+$G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u.$ (77)
+Assume that
+$\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0.$ (78)
+Then,
+$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (79)
+Proof Since $X$ has density $f$ and takes values in $0,\infty)$,
+$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t.$ (80)
+The survival function satisfies
+$\begin{aligned} \end{aligned}S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f\left(u\right)\mathrm{d}u,$ (81)
+so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$. Hence,
+$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t.$ (82)
 We now integrate by parts, using $G^{\prime}\left(t\right)=g\left(t\right)$,
 $-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t=-\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}+\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (83)
-By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$.
-Therefore, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
-A.1 Gaussian Distribution
+By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$. Therefore,
+$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (84)
+A.5.1 Gaussian Distribution
 Definition of sub-Gaussian.
-A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if $\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
+A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if
+$\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ (85)
+for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
 Basic properties of sub-Gaussian distributions.
@@
 For a Gaussian distribution, the optimal variance proxy is the same as the variance.
-Chernoff bound on the tail: $\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ for all $t\geq 0$.
+Chernoff bound on the tail:
+$\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ (86)
+for all $t\geq 0$.
 Maximum inequality.
-Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$.
-Denote $\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ as the smallest variance proxy for $X_{1},\dots,X_{d}$.
-Using Lemma 5,
-$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (27)
+Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$. Denote
+$\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ (87)
+as the smallest variance proxy for $X_{1},\dots,X_{d}$. Using Lemma 4,
+$\mathbb{E}\max_{k}X_{k}^{2}=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t$ (88)
+$=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$
 $\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
 $\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
-$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
+$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t$
+$=$ $\ \left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
 $=$ $\ t_{\star}^{2}+4d\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t_{\star}^{2}\right).$
@@
 $\mathbb{E}\max_{k}X_{k}^{2}\leq 2\sigma_{\star}^{2}\ln\left(2d\right)+2\sigma_{\star}^{2}=\left(2\ln\left(2d\right)+2\right)\max_{k}\mathbb{E}X_{k}^{2}$ (89)
-without needing the independence of $X_{k}$.
-Also, considering the trivial bound $\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2}$, then $\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}$.
-A.2 Laplacian Distribution
+without needing the independence of $X_{k}$. Also, considering the trivial bound
+$\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2},$ (90)
+then,
+$\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}.$ (91)
+A.5.2 Laplacian Distribution
 For a random variable $X\sim\mathrm{Laplace}\left(\mu,b\right)$ and $t\geq 0$,
@@
 The variance $\mathbb{E}\left(X-\mathbb{E}X\right)^{2}=2b^{2}$.
-Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$.
-Denote $b_{\star}=\max_{k}b_{k}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}$.
-For $t\geq 0$, $\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right)$.
+Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$. Denote
+$b_{\star}=\max_{k}b_{k}=\max_{k}\sqrt{\frac{1}{2}\mathbb{E}X_{k}^{2}}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}.$ (93)
+For $t\geq 0$,
+$\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right).$ (94)
 Using Lemma 4,
-$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (30)
+$\mathbb{E}\max_{k}X_{k}^{2}=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t$ (95)
+$=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$
 $\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
 $\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
-$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
+$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t$
+$=$ $\ \left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
 $=$ $\ t_{\star}^{2}+2db_{\star}\left(t_{\star}+b_{\star}\right)\exp\left(-b_{\star}^{-1}t_{\star}\right).$
-Splitting at $t_{\star}=b\ln d$ gives
+Splitting at $t_{\star}=b_{\star}\ln d$ gives
 $\mathbb{E}\max_{k}X_{k}^{2}\leq b_{\star}^{2}\left(\left(\ln d\right)^{2}+2\ln d+2\right)=\left(2^{-1}\left(\ln d\right)^{2}+\ln d+1\right)\max_{k}\mathbb{E}X_{k}^{2}$ (96)
 without needing the independence of $X_{k}$.
+Appendix B Further Experimental Results
+In this section, we report additional evaluations conducted across the Llama 3 and Qwen 3 family models in Tables˜4, 8, 3, 9, 6 and 7.
+For the Qwen models, the thinking mode is disabled across all experiments.
+The evaluation results are obtained using emulated (“fake”) quantization, where quantized values are represented in the BF16/FP16 format.
+Results are aggregated over multiple random seeds.
+Table 3: KL-divergence for Qwen3-8B measured on a slice of C4 dataset, with calibration performed on FineWeb dataset.
+Results (lower is better) are averaged over multiple random seeds.
+H denotes the Hadamard transform.
+Model Format Quantization Transform KL
+Qwen3-8B NVFP4 GPTQ H 0.069518
+RTN WUSH 0.065747
+GPTQ WUSH 0.054646
+MXFP4 GPTQ H 0.093119
+RTN WUSH 0.093739
+GPTQ WUSH 0.077909
+Table 4: W4A4 accuracy results on the LM Evaluation Harness for different models with RTN, GPTQ quantizations, and identity (I), Hadamard (H), WUSH transforms.
+For the Llama model, the MMLU-CoT metric is reported in the MMLU column.
+Model Format Quantization Transform MMLU GSM8K HellaSwag WinoGrande Average Recovery
+Llama-3.2-3B-Instruct BF16 - - 64.43 78.01 73.42 70.09 71.49 100.0
+NVFP4 RTN I 61.07 72.01 70.90 66.77 67.69 94.68
+H 59.91 64.82 69.77 65.59 65.02 90.95
+WUSH 59.91 71.24 70.96 67.25 67.34 94.19
+GPTQ H 60.22 70.91 71.06 66.57 67.19 93.99
+WUSH 60.17 72.71 71.26 67.32 67.87 94.93
+MXFP4 RTN I 56.81 60.80 67.30 64.56 62.37 87.24
+H 55.99 61.11 68.36 64.09 62.39 87.27
+WUSH 59.54 69.49 70.20 66.47 66.43 92.92
+GPTQ H 60.24 68.84 69.58 66.08 66.18 92.58
+WUSH 59.39 68.26 69.88 67.07 66.15 92.53
+Qwen3-8B BF16 - - 72.98 90.90 75.52 70.56 77.49 100.0
+NVFP4 RTN I 70.47 88.40 74.44 69.53 75.71 97.70
+H 70.19 86.35 73.02 68.11 74.42 96.04
+WUSH 70.86 89.92 74.84 70.02 76.42 98.61
+GPTQ H 70.78 89.11 74.34 69.25 75.87 97.91
+WUSH 71.03 89.56 74.88 70.32 76.45 98.66
+MXFP4 RTN I 67.69 84.23 71.24 67.40 72.64 93.74
+H 67.53 87.04 71.48 67.64 73.42 94.75
+WUSH 69.96 86.90 73.87 68.92 74.91 96.68
+GPTQ H 69.58 88.28 73.75 69.25 75.22 97.07
+WUSH 70.40 88.80 74.36 69.19 75.69 97.68
+Qwen3-14B BF16 - - 77.18 91.96 79.84 74.27 80.81 100.0
+NVFP4 RTN I 75.20 90.22 77.38 73.32 79.03 97.80
+H 74.98 92.04 77.76 72.38 79.29 98.12
+WUSH 75.43 91.78 78.82 73.18 79.80 98.76
+MXFP4 RTN I 72.92 90.22 76.68 71.51 77.83 96.31
+H 73.13 87.41 76.29 71.67 77.13 95.44
+WUSH 74.56 91.87 78.18 73.15 79.44 98.30
+Qwen3-32B BF16 - - 80.74 92.12 83.96 76.95 83.44 100.0
+NVFP4 RTN I 78.92 90.14 82.60 76.64 82.08 98.37
+H 79.48 91.81 82.51 74.98 82.20 98.51
+WUSH 79.56 92.72 83.14 75.79 82.80 99.24
+MXFP4 RTN I 76.82 75.28 81.32 74.74 77.04 92.33
+H 78.63 92.57 81.87 75.53 82.15 98.45
+WUSH 78.70 90.83 82.63 75.87 82.01 98.28
+Figure 6: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks.
+The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
+Figure 7: Comparison of different transforms on Qwen3-32B for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks.
+The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
+Figure 8: Comparison of different transforms on Llama-3.2-3B-Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks.
+The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
+Figure 9: Comparison of different transforms on Llama-3.1-8B-Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks.
+The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.